### 19 Sep 2024 v2

# Excess cataloged X-ray and radio sources at galaxy-cluster virial shocks

Abstract
We detect a highly significant excess of X-ray (2RXS) and radio (NVSS, GMRT, VLSSr) catalog sources when stacked around MCXC galaxy clusters and groups, narrowly confined within $\lesssim 100\mathrm{\,kpc}$ of the $\sim 2.4R_{500}$ virial shock radius (inferred from previous continuum stacking), with similar X-ray ($\sim 4\sigma$ for $443$ clusters) and radio ($\sim 4\sigma$ for $485$ clusters) characteristics ($>5\sigma$ joint). The excess sources show 10101010100100100100 kpc scales, $L_{X}(0.1\mbox{--}2.4\mbox{ keV})\simeq 10^{42\mbox{\scriptsize--}43}\mathrm{\,erg\,s^{-1}}$ or $\nu L_{\nu}(\nu=1.4\mathrm{\,GHz})\simeq 10^{40{\mbox{\scriptsize--}}41}\mathrm{\,erg\,s^{-1}}$ luminosities, and a preferentially radial radio-polarization. The narrow localization and properties of the excess identify these sources not as AGN, often invoked speculatively for excess X-ray sources at cluster outskirts, but rather as infalling gaseous clumps interacting with the virial shock, probably galactic halos and possibly outflow remnants. The local excess of such discrete, radio-to-$\gamma$-ray sources around an object can probe its virial shock also at high redshifts and sub-cluster scales.
1 Introduction
Several studies have reported a radially-broad excess of X-ray sources in the outskirts of galaxy clusters. These include a $\gtrsim 5\sigma$ excess at $1.5\lesssim r\lesssim 3\mbox{ Mpc}$ radii in stacked Chandra data around 24 relaxed, redshift $0.70.3<z<$ 0.7 MACS clusters [ruderman2005origin], a $\sim 4\sigma$ excess at $2\lesssim r\lesssim 3\mbox{ Mpc}$ in stacked XMM-Newton data around 22 luminous, $0.9<z\lesssim 1.6$ clusters [FassbenderEtAl12], and a statistically significant excess at normalized, $42<\tau\equiv r/R_{500}<$ 4 radii in stacked XMM-LSS field data around 19 clusters of $0.350.14<z<$ 0.35 [koulouridis2014x]. Here, the subscript $500$ refers (henceforth) to the radius around the center of a cluster enclosing a mass density $\rho$ that is $500$ times larger than the critical mass density of the Universe. The excess sources were usually assumed to be associated with active galactic nuclei (AGN). Their excess at the cluster periphery was tentatively attributed to an enhanced rate of major galaxy mergers [e.g., ruderman2005origin, EhlerEtAl15, and references therein]. Part of the excess was speculated to arise from gravitational lensing of background quasi-stellar objects [koulouridis2014x] or from foreground structures [FassbenderEtAl12].
In contrast, other analyses, including larger studies, have found no excess of peripheral sources. In particular, no X-ray source excess was found beyond $r=1\mbox{ Mpc}$ using Chandra data around 27 disturbed, $0.70.3<z<$ 0.7 MACS clusters [ruderman2005origin] or 148 clusters of $0.90.1<z<$ 0.9 [GilmourEtAl09], and no excess of optical, IR or radio-selected cataloged AGN was found at the outskirts of 2300 infrared-selected clusters [mo2018massive]. Searches for a source excess performed after normalizing distances to typical cluster scales such as $R_{500}$ did not show a consistent signal, either; for example, a local deficit of X-ray sources was found at $32<\tau<$ 3 around 14 clusters of $1.050.43<z<$ 1.05 [koulouridis2014x].
More recently, a $\sim 3\sigma$ excess of Chandra X-ray sources was reported [koulouridis2019high] in a fairly narrow, $2.52<\tau<$ 2.5 range of normalized radii around a small sample of five $z\sim 1$, mostly disturbed clusters. This signal is dominated by two clusters, each with about ten excess $2.52.0<\tau<$ 2.5 sources with respect to neighbouring bins. The excess sources show $10^{42.5}{\rm{\,erg\,s^{-1}}}\lesssim L_{0.5\mbox{ \scriptsize keV}}^{8.0\mbox{ \scriptsize keV}}\lesssim 10^{44}{\rm{\,erg\,s^{-1}}}$ luminosities, where energy or frequency subscripts (superscripts) denote low (high) band limits, henceforth. These sources were again tentatively attributed to enhanced AGN activity associated with excessive galaxy mergers or gas stripping.
However, an excess of AGN over such a narrow range of radii at the outskirts of clusters would be unnatural, especially taking into account projection (and, in other studies, also stacking) effects. Indeed, the rate of galaxy mergers is thought to evolve quite gradually, rising well before the accreted galaxies reach the cluster outskirts, and remaining high after pericentric crossing [e.g., VijayaraghavanRicker13]. AGN activity can be affected by the intracluster medium (ICM) through several channels, notably, in this context, by ram pressure stripping of hot halo gas [e.g., McCarthyEtAl08], but these processes are thought to be important only once the galaxy approaches the central, $r\lesssim 1\mbox{ Mpc}$ region of the cluster [VijayaraghavanRicker13, and references therein].
Therefore, a sharp spike in the radial profile of the density of sources would indicate a localized, sudden change in the environment. Interestingly, the structure formation, or virial, shock of a galaxy cluster, long anticipated to lie in the $32<\tau<$ 3 range according to cosmological simulations [e.g., KeshetEtAl03, SchaalSpringel15], provides a natural candidate for such a sharp environmental transition. Indeed, as an infalling object crosses the virial shock (VS), it should see the ambient pressure suddenly jump by orders of magnitude as the infalling surrounding plasma is violently slowed down and heated, accompanied by the generation of magnetic fields and shock-accelerated relativistic particles (cosmic rays, CRs) carrying a substantial fraction of the elevated thermal energy. While crossing a VS is unlikely to trigger an AGN, and the processes by which it could lead to the emergence of bright excess X-ray or radio sources may be a-priori unclear, the VS could explain at least the position and sharpness of the signal.
Moreover, recent studies have identified stacked VS signals in precisely the same $2.52<\tau<$ 2.5 region harboring the excess [koulouridis2019high] X-ray sources. Such stacking analyses include a $>5\sigma$ signal at $2.2\lesssim\tau\lesssim 2.5$ in Fermi-LAT $\gamma$-ray data around 112 extended clusters, interpreted as CR electrons (CREs) inverse-Compton (IC) scattering cosmic microwave background (CMB) photons [ReissEtAl17, reiss2018detection]; a $\gtrsim 5\sigma$ synchrotron counterpart at $2.4\lesssim\tau\lesssim 2.6$ in OVRO-LWA radio data around 44 massive clusters [HouEtAl23]; a significant drop around $2.0\lesssim\tau\lesssim 2.6$ in Planck $y$-parameter around 10 galaxy groups [PrattEtAl21]; and a projected $y$-parameter drop starting at $\tau\simeq 2$ in South Pole Telescope data around 500 clusters [Anbajagane2022]. Such stacked VS signals corroborate the IC, synchrotron, and Sunyaev-Zel’dovich (SZ) signals detected separately or simultaneously in individual nearby clusters [KeshetEtAl17, keshet2018evidence, HurierEtAl19, keshet20coincident], and roughly coincide with the splashback feature inferred from a localized drop in the logarithmic slopes of galaxy density profiles [MoreEtAl16, ShinEtAl19]. Interestingly, stacked data suggest a secondary, weaker and broader IC [reiss2018detection], synchrotron [HouEtAl23], and possibly also SZ [Anbajagane2022] signal around $\tau\simeq 6$; if verified, this would suggest an elongated averaged VS morphology of $\sim 2.5$ ellipticity ratio, as indicated in Coma [KeshetEtAl17, keshet2018evidence].
[…]
The paper is organized as follows. We describe the cluster and source catalogs, along with our selection criteria, in §2, and outline our stacking and modelling methods in §3. The results are presented and phenomenologically modelled in §4. The properties of the excess sources are inferred statistically in §5, and their physical origin is explored in §6. Finally, our results are summarized and discussed in §7.
The Supplementary material provides supporting and auxiliary information regarding the spectra of the excess X-ray sources (§A), limitations on source size estimates (§B), the evaluation of excess confidence levels (CLs) under Poisson statistics (§C), various convergence and sensitivity tests (§D), additional properties of the excess sources (§E), and the absence of a significant VS signal from 2RXS–NVSS pairs (§F) or identified AGN (§G). All (field+excess) sources found near the VS are listed in §H.
We adopt a $\Lambda$CDM cosmological model with a Hubble constant $H_{0}=68\mbox{ km}\mbox{ s}^{-1}\mbox{ Mpc}^{-1}$, a mass fraction $\Omega_{m}=0.3$, and a $76\%$ hydrogen mass fraction, yielding a mean particle mass $\bar{m}\simeq 0.59m_{p}$. The center of a galaxy cluster is defined as its X-ray peak, and an adiabatic index $\Gamma=5/3$ is assumed for the ICM. Confidence intervals are $68\%$ containment projected for one parameter.
2 Data preparation
Here, we describe the catalogs and our data preparation procedure. The MCXC cluster catalog is discussed in §2.1, the 2RXS catalog of X-ray sources (subscript $X$) in §2.2, and the NVSS radio catalog (subscript $\nu$) in §2.3. Table 1 summarizes the main properties of the catalogs.
2.1 Cluster catalog and selection criteria
The MCXC catalog of galaxy clusters [PiffarettiEtAl11], based on a compilation of the ROSAT all-sky survey and ROSAT serendipitous catalogs, is particularly suitable for our purposes as it facilitates a uniform normalization to a fixed overdensity. The catalog, with 1743 nearby clusters, was used in previous stacking analyses [reiss2018detection, HouEtAl23], and provides the redshift and characteristic length scale $R_{500}$ of each cluster, so the corresponding angular scale $\theta_{500}$ can be estimated.
We apply the following nominal cuts to the MCXC, in order to select the clusters most suitable for source stacking and to avoid contaminations.
  1. 1.
    Sufficiently extended clusters to resolve the virial ring. The source catalog resolutions are $45{{}^{\prime\prime}}$ (FWHM in radio) and $\sim 2{{}^{\prime}}$ (detection cell in X-rays), whereas the thickness of the virial ring in continuum emission was estimated as $\sim\theta_{500}/3$ [reiss2018detection, HouEtAl23], so we conservatively require $\theta_{500}>3\cdot 2{{}^{\prime}}=6{{}^{\prime}}$. This cut is independently justified as it coincides with the MCXC turnaround angle, below which the number of detected clusters diminishes.
  2. 2.
    Clusters that are less extended than typical foreground structures. The power spectrum of sources in the 2RXS (NVSS) catalog shows substantial structure on $\gtrsim 3^{\circ}$ ($\gtrsim 1^{\circ}$) scales, suggesting extended foregrounds. To avoid contaminating the VS region with such putative structures, we impose an upper, $\theta_{500}\leq 0{\overset{{}^{\circ}}{.}}3$ limit.
  3. 3.
    Sufficiently massive galaxy groups. In order to focus on galaxy clusters and rich galaxy groups, but avoid individual galaxies and poor groups which may not show the same type of peripheral sources, we impose a lower $M_{500}\geq 10^{13}{M_{\odot}}$ mass limit.
  4. 4.
    Avoiding confusion in regions of a high sky-projected density of clusters. The MCXC shows such high densities, with nearly overlapping projected virial radii, in two regions, around Galactic coordinates $\{l,b\}\sim\{315^{\circ},32^{\circ}\}$ and $\sim\{12^{\circ},50^{\circ}\}$ [reiss2018detection]. We avoid these regions, by excluding all clusters within $5^{\circ}$ of these two sky positions. This cut is in practice redundant, as the other cuts already exclude all clusters in these two busy regions.
  5. 5.
    Avoiding large Galactic structures. The Loop-I and Fermi bubble regions are excluded from the nominal analysis, to avoid any possible effects they may have on the catalogs. For this purpose, we adopt simplified models of these two structures. Loop-I is modeled as a disk centered on $\left\{l,b\right\}=\left\{329,17.5\right\}$, with a radius of $55^{\circ}$ [e.g. Haslam1970, Dickinson2018]. The Fermi-bubble is crudely modeled as the region defined by $\left|b\right|\leq 55^{\circ}$ and either $l<30^{\circ}$ or $l>330^{\circ}$.
  6. 6.
    Additional cuts are imposed on the Galactic latitude $|b|$, to avoid the bright Galactic plane, and on the declination, Dec. These cuts are specific to each source catalog, so are detailed below.
The above cuts leave a reduced sample of ${N}_{c,X}=443$ clusters for stacking X-ray sources, and a sample of ${N}_{c,\nu}=485$ clusters for radio stacking. These clusters have redshifts in the range $0.240.017<z<$ 0.24, and masses in the range $10^{13}{M_{\odot}}<M_{500}<9\times 10^{14}{M_{\odot}}$; these and additional sample properties are provided in Table 1. The MCXC includes $\sim 50$ ($\sim 90$) clusters which harbour another MCXC cluster inside (twice) their virial radius, about $80\%$$90\%$ of which survive our cluster cuts. These clusters with nearby neighbours are not excluded from our nominal analysis, but we verify in §D that excluding close pairs does not alter our results significantly.
2.2 X-ray source catalog and cuts
[…]
Softer X-rays undergo more photo-absorption by neutral hydrogen atoms in the Galaxy. The resulting optical depth exceeds unity for $\epsilon\simeq 0.1\mbox{ keV}$ photons at $\left|b\right|\lesssim 30^{\circ}$ Galactic latitudes. Therefore, for the nominal X-ray analysis, we place an extended cut on the Galactic latitude of the clusters, retaining only those with $\left|b\right|\geq 30^{\circ}$.
The number of sources in a catalog diminishes as one increases the threshold on the minimal likelihood $\mathcal{L}$ of source detection, defined such that the false-positive source probability is $e^{-\mathcal{L}}$. The decline in the fraction of spurious sources as $\mathcal{L}$ increases has been estimated using simulations [2RXS_paper]. The 2RXS catalog compromises on $\mathcal{L}\geq 6.5$, corresponding to $\gtrsim 3\sigma$ source detection with a $\sim 30\%$ fraction of spurious sources in the catalog. Aiming for a smaller, $\sim 10\%$ fraction of spurious sources, we adopt $\mathcal{L}\geq 8$ in our nominal analysis, corresponding to $\gtrsim 3.4\sigma$ source detection.
Figure 1: Hammer-Aitoff all-sky projection of source density (cubehelix [Green11_Cubehelix] color map) and cluster sample (orange circles of radius $5{\theta_{500}}$), for X-ray (left panel; 2RXS) and radio (right panel, NVSS; note diminished exposure at low declinations, with no $\mbox{Dec}<-40^{\circ}$ sources) analyses. Also shown are the cuts of the Galactic plane (dotted yellow lines), the Fermi bubbles (dashed red curves), and Loop-I (dot-dashed cyan).
Table 1: Catalog and stacking properties.
Property Symbol 2RXS NVSS
[…]
Sky coverage All-sky $\mbox{Dec}>-40^{\circ}$ ($82\%$ sky)
[…]
Detection threshold $\mathcal{L}>6.5$ $I_{\nu}>2\mbox{ mJy}\rm{\,beam^{-1}}$
Base map resolution $\sim 45{{}^{\prime\prime}}$ (pixel size) $45{{}^{\prime\prime}}$ (FWHM)
Source size (see §B) $r,\{a,b\}$ $r\geq r_{psf}\simeq 40^{\prime\prime}$ $14^{\prime\prime}<2b<2a<286^{\prime\prime}$
Total source number $1.3\times 10^{5}$ $1.8\times 10^{6}$
Cuts Latitude cut $b$ $|b|>30^{\circ}$ $|b|>10^{\circ}$
Cluster declination cut Dec $\mbox{Dec}>-35^{\circ}$
Source fluxa cut $F^{*}$ $F_{X}<F_{X}^{*}\equiv 7\times 10^{-12}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ $F_{\nu}>F_{\nu}^{*}\equiv 100\mbox{ mJy}$
Source likelihood cut $\mathcal{L}$ $\mathcal{L}>8$
Clustersb Number of clusters ${N}_{c}$ ${N}_{c,X}=443$ ${N}_{c,\nu}=485$
[…]
Angular radius median (range) $\theta_{500}$ $150{\overset{{}^{\circ}}{.}}$ 15 ($300{\overset{{}^{\circ}}{.}}10{\mbox{--}}0{\overset{{}^{\circ}}{.}}$ 30) $140{\overset{{}^{\circ}}{.}}$ 14 ($300{\overset{{}^{\circ}}{.}}10{\mbox{--}}0{\overset{{}^{\circ}}{.}}$ 30)
[…]
Resultsb Mean field source density $f$ 5360 $5360\mbox{ sr}^{-1}$ 5760 $5760\mbox{ sr}^{-1}$
ICM-region sources (expected field) $\mathcal{N}$ ($\mathcal{F}$) 107 ($\sim$82) 102 ($\sim$92)
VS-region sources (expected field) $\mathcal{N}$ ($\mathcal{F}$) 119 ($\sim$77) 128 ($\sim$87)
VS SW-overdensity (significance) $\delta_{\rm SW}$ $0.54$ ($4.4\sigma$ Poisson-corrected) $0.47$ ($4.0\sigma$)
  • a — Energy flux (specific energy flux) in X-rays (radio). The cuts are explained in the text.
    b — After cuts. Source numbers $\mathcal{N}$ are summed over all clusters. VS sources are listed in §H.
We define the X-ray energy flux $F_{X}\equiv F_{0.1\mbox{ \scriptsize keV}}^{2.4\mbox{ \scriptsize keV}}$ as the specific flux $F_{\epsilon}$ integrated over the ROSAT band. The catalog provides flux estimates based on three different emission models: a power-law spectrum, the optically-thin thermal plasma emission model Mekal [MeweEtAl86], and optically thick, black-body emission. In most or all relevant cases, the model fits are not sufficiently accurate to identify which emission model is most appropriate. Therefore, we henceforth adopt the power-law estimate of $F_{X}$, which is a simpler fit available for the largest number of sources; our results are not sensitive to this choice (see §D). We nevertheless consider below the fitted source temperatures, as well as the fitted photon power-law index $\Gamma_{X}$ defined by $F_{\epsilon}\propto\epsilon^{1+\Gamma_{X}}$.
Figure 2 (left panel) shows the energy flux histogram of X-ray sources in the analyzed sky region. The 2RXS catalog is dominated by foreground sources brighter than the stacked clusters themselves. To reduce foreground confusion, we impose an upper flux limit that removes sources whose luminosities would exceed the typical, $L_{X}\simeq 10^{45}{\rm{\,erg\,s^{-1}}}$ luminosity of the entire ICM even around the most distant ($z_{\rm max}\simeq 0.23$) clusters in our sample, thus retaining only $F_{X}<F_{X}^{*}\equiv 7\times 10^{-12}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ sources. Note that this $F_{X}^{*}$ cutoff is already quite close to the 2RXS $\sim 10^{-13}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ sensitivity threshold [2RXS_paper]. In §D, we show that our results are not sensitive to the precise choice of flux cut; for example, a high $F_{X}^{*}=10^{-10}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ changes the significance of excess virial sources by only $\sim 2\%$. In fact, a significant, $\sim 4\sigma$ signal persists even with no flux cut at all, despite the strong foreground.
Figure 2: The logarithmic energy flux $F_{X}$ (specific flux $F_{\nu}$) histogram of 2RXS (NVSS) sources in the analyzed sky is shown in the left (right) panel, along with our nominal flux thresholds (black vertical lines with explanatory labels). Our results are not sensitive to the precise choice of thresholds; see text.
2.3 Radio source catalog and cuts
We use the extensive NVSS catalog for radio sources, unless otherwise stated. This catalog covers only $\sim 82\%$ of the sky, but radio waves are not substantially absorbed in the galaxy, so the catalog can be used at lower latitudes than in X-rays. Consequently, NVSS allows for the stacking of a similar and even slightly larger number of clusters than with 2RXS. The projected density $n_{\nu}$ of catalog radio sources on the sky is shown in Fig. 1 (right panel). Table 1 summarizes some relevant properties of the catalog and of its sources (in column 4).
The NVSS catalog covers the northern sky at declinations $\mbox{Dec}>-40^{\circ}$ (J2000). Figure 1 shows an apparent decline in $n_{\nu}$ with decreasing Dec even somewhat north of this declination. Therefore, we impose a corresponding cut on the declination of clusters used for radio stacking, nominally keeping only those with $\mbox{Dec}>-35^{\circ}$.
[…]
A histogram of the specific flux of NVSS sources is shown in the right panel of figure 2. This catalog is strongly dominated by faint background sources (in contrast to the bright foreground source of 2RXS), due in part to star-forming galaxies, whose luminosity function peaks near $\nu L_{\nu}\sim 10^{39}{\rm{\,erg\,s^{-1}}}$ [e.g., radio_Lum]; for the most part, these sources are not expected to be associated with the low-redshift clusters in our sample. We therefore impose a lower specific-flux limit that removes sources with luminosities lower than this peak luminosity even around the nearest ($z_{\rm min}\simeq 0.017$) clusters in our sample, thus retaining only sources of $F_{\nu}\gtrsim F_{\nu}^{*}\equiv 100\mbox{ mJy}$. In §D, we show that changing the radio flux cutoff by $50\%$ has only a $\sim 15\%$ effect on the significance of our results, and our choice is conservative. Interestingly, the signal persists even if one avoids a flux cut entirely, thus retaining the large number of faint background sources, but expectedly diminishes to the $\sim 2\sigma$ level.
3 Stacking analysis method
We normalize sky distances around each cluster by its $\theta_{500}$, radially bin the data onto circular annuli about the center of the cluster, and stack the resulting normalized, binned, radial distribution over all clusters in the sample. Stacking is performed for X-ray and radio sources separately, although we later consider also the joint distribution and the significance of the joint excess. The following procedure thus pertains separately to the stacking of 2RXS and NVSS sources, as well as GMRT and VLSSr sources considered later.
For each cluster, we define the normalized radius $\tau\equiv\theta/{\theta_{500}}$, where $\theta$ is the angular distance from the center of the cluster. Alternatively, in order to generate stacked 2D maps (shown in §4 and in §D), i.e. without radial binning, we replace $\tau$ by its two-dimensional analog ${\bm{\tau}}\equiv\{\tau_{x},\tau_{y}\}$. For the purpose of stacking, we define the region of interest (ROI) as $\tau<{\tau_{\rm max}}=10$, choosing ${\tau_{\rm max}}$ to be about four times larger than the VS radius inferred from stacking Fermi-LAT or LWA data. Our results are not sensitive to the choice of ${\tau_{\rm max}}$ and all other stacking parameters, as shown in §D.
In order to test for a putative local excess of sources and determine its significance, one must estimate the sky-projected density of background or foreground sources, henceforth referred to colloquially as field sources, in the vicinity of each cluster. In our nominal approach, we consider a fixed projected density of field sources around each cluster, measured outside its ROI, in the so-called field region defined by ${\theta_{\rm min}}<\theta<{\theta_{\rm max}}$. We choose ${\theta_{\rm min}}=3^{\circ}$ slightly larger than the ROI radius, i.e. ${\theta_{\rm min}}>{\tau_{\rm max}}\theta_{500}$, of any cluster in our sample. We take ${\theta_{\rm max}}=5^{\circ}$ sufficiently larger than ${\theta_{\rm min}}$, for good statistics and to overcome foreground fluctuations over $\gtrsim 1^{\circ}$ scales. Alternative choices of $[{\theta_{\rm min}},{\theta_{\rm max}}]$, as well as more elaborate, non-constant field models, are examined in §D.
When testing for virial-shock signals, it can be useful to examine, in addition to the field region outside the VS, also the region just inward of the shock, beyond its range of influence. For this purpose, we designate the range $1.51.0<\tau<$ 1.5 as a so-called ICM region. The upper limit on this region is chosen sufficiently far from the virial radius to reduce the projected effects of the shock. The lower limit is chosen to balance sufficient statistics with limited evolutionary effects and central-source contamination.
[…]
For each catalog, $\tau$ bin, and cluster ${\mathsf{c}}$, we compute the solid angle $\Delta\Omega(\tau,{\mathsf{c}})$ and number $\mathcal{N}(\tau,{\mathsf{c}})$ of sources falling within the bin. The dimensionless projected density of catalog sources can now be estimated as
$n(\tau,{\mathsf{c}})=\frac{\mathcal{N}(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}\mbox{ ,}$ (3.1)
where
$\Delta\omega(\tau,{\mathsf{c}})\equiv\frac{\Delta\Omega(\tau,{\mathsf{c}})}{\theta_{500,{\mathsf{c}}}^{2}}$ (3.2)
is the solid angle normalized to a square $\theta_{500}^{2}$ pixel. The number $\mathcal{F}(\tau,{\mathsf{c}})$ of field sources anticipated in the bin is extrapolated from the field region, giving the analogous dimensionless projected field density
$f(\tau,{\mathsf{c}})=\frac{\mathcal{F}(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}\mbox{ ,}$ (3.3)
and the implied fractional overdensity
$\delta(\tau,{\mathsf{c}})\equiv\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})}{\mathcal{F}(\tau,{\mathsf{c}})}=\frac{n(\tau,{\mathsf{c}})-f(\tau,{\mathsf{c}})}{f(\tau,{\mathsf{c}})}\mbox{ .}$ (3.4)
In our nominal, uniform field model, with a constant dimensionless projected density $f({\mathsf{c}})$ around each cluster ${\mathsf{c}}$, the field distribution becomes $\mathcal{F}(\tau,{\mathsf{c}})=f({\mathsf{c}})\Delta\omega(\tau,{\mathsf{c}})$.
In the $\mathcal{F}\gg 1$ limit, one may approximate its distribution as normal, as we verify using control cluster samples in §3.3. We may then infer the Z-test significance of any $\mathcal{N}>\mathcal{F}$ excess with respect to the field-only null hypothesis,
$S\simeq\frac{\mathcal{N}-\mathcal{F}}{\sqrt{\mathcal{F}\,}}\mbox{ .}$ (3.5)
In practice, individual cluster bins at small radii typically show $\mathcal{F}<1$, so the significance $S$ should be estimated from Poisson statistics (see §C) or from control clusters (see §3.3). One may co-add the $N_{\mbox{\tiny{cat}}}=2$ catalogs, similarly estimating the joint Z-test significance
$S_{\mbox{\tiny{joint}}}\simeq\frac{1}{\sqrt{N_{\mbox{\tiny{cat}}}}}\sum_{i}^{N_{\mbox{\tiny{cat}}}}S_{i},$ (3.6)
where $S_{i}$ is the significance profile in each catalog.
There is some freedom in the method in which these quantities may be stacked over clusters. Following previous MCXC stacking analyses [reiss2018detection, HouEtAl23], we consider two extremes. In the so-called source-weighted (SW) co-addition method, discussed in §3.1, direct quantities (source numbers, solid angles) are stacked separately over all clusters and then combined to infer derived quantities (source densities, overdensities, excess significance levels). In the cluster-weighted (CW) co-addition method, discussed in §3.2, derived quantities are first computed on a cluster-by-cluster basis, and only subsequently stacked. The results in §4 show good agreement between the two methods in the VS region, indicating that many clusters contribute fairly evenly to the stacked signal.
The nominal significance levels $S$ obtained above and below, based on approximating the distribution as normal and using standard error propagation, are generalized for Poisson statistics (see §C), and tested using samples of control clusters, as described in §3.3. We present simplified models for the radial dependence of the excess in §3.4, and discuss in §3.5 the procedure of fitting these models to the data and assessing them using the TS-test.
3.1 Source-weighted co-addition
The SW stacking method assigns an equal weight to each source, rather than to each cluster. In each normalized angular $\tau$ bin, we thus separately co-add, over all clusters, both the number of sources,
$\mathcal{N}(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\mathcal{N}(\tau,{\mathsf{c}})\mbox{ ,}$ (3.7)
and the solid angle
$\Omega(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\Delta\Omega(\tau,{\mathsf{c}})\mbox{ ,}$ (3.8)
where ${N}_{c}$ is the number of clusters in the sample.
The dimensionless projected number density of sources, co-added over clusters in the bin $\tau$, can now be computed as
$n_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{N}(\tau)}{\bar{\omega}(\tau)}\mbox{ ,}$ (3.9)
where the dimensionless $\bar{\omega}(\tau)\equiv\Omega(\tau)/\bar{\theta}_{500}^{2}$ is obtained by normalizing to the mean $\bar{\theta}_{500}^{2}\equiv{N}_{c}^{-1}\sum_{{\mathsf{c}}}\theta_{500}$; this normalization has no effect on our subsequent results. The statistical uncertainty in $n_{\rm\mbox{\tiny{SW}}}$ may be estimated as
$\sigma[n_{\rm\mbox{\tiny{SW}}}(\tau)]\simeq\frac{\sqrt{\mathcal{N}(\tau)}}{\bar{\omega}(\tau)}\mbox{ .}$ (3.10)
Source-weighted estimates such as $n_{\rm\mbox{\tiny{SW}}}(\tau)$ are somewhat skewed toward extended clusters or clusters in strong-field regions, which may harbour a larger number of sources and so may dominate the numerator of Eq. (3.9). This bias is offset in part by the comparable extents of all clusters in our sample, and by avoiding regions with known contaminations.
The above approach is applied analogously to field sources, modelled with some distribution $\mathcal{F}(\tau,{\mathsf{c}})$ based on the field region. In analogy with Eqs. (3.7)–(3.9), the co-added number of field sources is
$\mathcal{F}(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\mathcal{F}(\tau,{\mathsf{c}})\mbox{ ,}$ (3.11)
with co-added dimensionless projected density
$f_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{F}(\tau)}{\bar{\omega}(\tau)}\mbox{ .}$ (3.12)
The significance of a local excess of sources with respect to the field-only null hypothesis can now be estimated as in Eq. (3.5),
$S_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{N}(\tau)-\mathcal{F}(\tau)}{\sqrt{\mathcal{F}(\tau)}}\mbox{ ,}$ (3.13)
in the $\mathcal{F}(\tau)\gg 1$ limit of a normal distribution.
Finally, the overdensity $\delta_{\rm\mbox{\tiny{SW}}}$ of sources with respect to the field can be estimated as
$\delta_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{n_{\rm\mbox{\tiny{SW}}}(\tau)}{f_{\rm\mbox{\tiny{SW}}}(\tau)}-1\mbox{ ,}$ (3.14)
with an associated uncertainty
$\sigma[\delta_{\rm\mbox{\tiny{SW}}}(\tau)]$ $\simeq$ $\frac{n_{\rm\mbox{\tiny{SW}}}(\tau)}{f_{\rm\mbox{\tiny{SW}}}(\tau)}\sqrt{\frac{\mathcal{N}(\tau)+\mathcal{F}(\tau)}{\mathcal{N}(\tau)\mathcal{F}(\tau)}}$
$\simeq$ $\left[1+\delta_{\rm\mbox{\tiny{SW}}}(\tau)\right]\sqrt{\frac{2}{f_{\rm\mbox{\tiny{SW}}}(\tau)\bar{\omega}(\tau)}}\mbox{ .}$
The second approximation holds when the excess is small compared to the field; it is shown here for completeness, but is not used in practice below.
3.2 Cluster-weighted co-addition
The CW stacking method assigns the same weight to each cluster, rather than to each source. In a given cluster ${\mathsf{c}}$, the significance of a local excess of sources with respect to the field-only null hypothesis is estimated as
$S(\tau,{\mathsf{c}})\simeq\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})}{\sqrt{\mathcal{F}(\tau,{\mathsf{c}})}}\mbox{ .}$ (3.16)
Under the null hypothesis, we may approximate $S(\tau,{\mathsf{c}})$ as a random variable of zero mean and unit standard deviation. The CW co-added significance of the excess over the field may now be estimated as
$S_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{\sqrt{{N}_{c}}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}S(\tau,{\mathsf{c}})\mbox{ .}$ (3.17)
Cluster-weighted estimates are somewhat skewed by clusters with low field values, where $|S(\tau,{\mathsf{c}})|$ can have artificially large values.
In the same method, we co-add densities over the sample of clusters to obtain the average stacked dimensionless projected density
$n_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum_{{\mathsf{c}}=1}^{{N}_{c}}n(\tau,{\mathsf{c}})\mbox{ ,}$ (3.18)
with the associated uncertainty
$\sigma[n_{\rm\mbox{\tiny{CW}}}(\tau)]=\frac{1}{{N}_{c}}\sqrt{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{n(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}}\mbox{ .}$ (3.19)
The co-added dimensionless projected density of field sources then become
$f_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}f(\tau,{\mathsf{c}})\mbox{ .}$ (3.20)
The local overdensity of sources above the field may be defined in each cluster as
$\delta(\tau,{\mathsf{c}})=\frac{n(\tau,{\mathsf{c}})}{f(\tau,{\mathsf{c}})}-1\mbox{ ,}$ (3.21)
so the stacked overdensity is
$\delta_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\delta\left(\tau,{\mathsf{c}}\right)\mbox{ ,}$ (3.22)
with an associated uncertainty
$\sigma\left[\delta_{\rm\mbox{\tiny{CW}}}(\tau)\right]=\frac{1}{{N}_{c}}\sqrt{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{n^{2}(\tau,{\mathsf{c}})}{f^{2}(\tau,{\mathsf{c}})\mathcal{N}(\tau,{\mathsf{c}})}}\mbox{ ,}$ (3.23)
where we neglected fluctuations in $\mathcal{F}$.
3.3 Poisson and control sample simulations
In the absence of any signals and source correlations, the statistical distribution of the small number of source counts in each radial bin within each cluster would follow a Poisson distribution, with a source density fixed by the foreground. However, the statistical significance and uncertainty estimates in Eqs. (3.5)–(3.6), (3.10), (3.1)– (3.17), (3.19), and (3.23) invoke, for simplicity, normally-distributed foreground counts. We run Monte-Carlo Poisson simulations to test and correct for this non-Poissonian approximation. Such simulations do not, however, capture structures and correlations in the actual catalogs. We therefore study, in addition, the statistics of the actual catalog sources around a large number of control (or mock) cluster samples, in order to test for systematics and validate our method.
Our Monte-Carlo Poisson simulations involve drawing, from a Poisson distribution, $10^{6}$ random foreground source-counts for each $\tau$ bin in each cluster, based on the estimated foreground source-density $\mathcal{F}(\tau,{\mathsf{c}})$ of the cluster in each catalog. These counts represent $\mathcal{N}(\tau,{\mathsf{c}})$ under the field-only null hypothesis.
The control sample analysis is based on $\sim 10^{6}$ samples of control clusters for each catalog, each such sample including ${N}_{c}$ clusters with the same ${\theta_{500}}$ values as in the real cluster sample, but placed in random positions within the allowed, i.e. non-masked, part of the sky. We stack the real catalog sources around each such control cluster, and calculate the corresponding distributions of foreground and excess source properties, just as in the real sample.
For both Monte-Carlo simulations and control-sample tests, we then study the resulting distributions of $S(\tau)$ values, each quantifying the significance of the putative excess above the foreground in a particular instance. Figure 3 shows the ±$\pm 1\sigma,\pm 2\sigma,\ldots$ CLs of $S$ around the median of its distribution, as inferred from the corresponding $68\%,95\%,\ldots$ containment values. These CLs are shown based on both Poisson simulations and control samples, for both X-ray and radio catalogs, using both SW and CW stacking methods.
As expected, the results approach Eqs. (3.13) and (3.17) at large $\tau$, where $\mathcal{F}\gg 1$, but deviate substantially at small $\tau$. The deviations are particularly large for $\mathcal{N}<\mathcal{F}$ source deficits, but we are mostly interested in the $S>0$ significance of a positive excess. The approximate agreement of the $S=0,+1,+2\ldots$ curves with the corresponding Poisson and control-sample containment fractions confirms that our nominal §3.1 and §3.2 estimates capture the significance of the excess reasonably well, with small errors $\Delta S$ in $S$. Quantitatively, for the relevant, $\tau>1$ range, we find modest, $|\Delta S|/S\lesssim 10\%$ fractional errors, except for radio CW stacking where $|\Delta S|/S\lesssim 15\%$. While small, this $\Delta S$ should be accounted for at the VS region.
Much of the $\Delta S$ difference is attributed to the inaccurate normal-distribution approximation of Poisson statistics. Indeed, the Poisson and control-sample estimates are in much better agreement with each other, for both positive and negative $S$, and even for $\tau<1$. The difference $\Delta S$ between them is very small in X-rays, where $|\Delta S|/S\simeq 4\%$ ($|\Delta S|/S\simeq 2\%$) for $S>0$ ($S<0$), and is acceptable for radio SW stacking, where $|\Delta S|/S\simeq 6\%$ ($|\Delta S|/S\simeq 2\%$). For these cases, we henceforth apply an analytic correction (§C), correcting our nominal-distribution estimates for Poisson statistics; the relevant Eqs. (C.1)–(C.2) are shown to accurately follow the simulated Poisson results in Fig. 3. The overall agreement between Poisson and control-sample results indicates that systematic effects and correlations are weak, and our Poisson-statistics estimates are reliable.
For radio CW stacking, however, Poisson and control-sample estimates differ by a more substantial $\Delta S/S\simeq 10\%$ ($\Delta S/S\simeq 3\%$) for $S>0$ ($S<0$), perhaps due to extended sources or sidelobes misidentified as multiple sources. In this case, we henceforth interpolate the control-sample CLs in order to obtain a more accurate estimate of $S$.
Figure 3: Statistical foreground analysis of SW (left panels) and CW (right) stacked X-ray (upper panels) and radio (bottom) sources. The ±$\pm 1,\pm 2,\pm 3,\ldots$ CLs of control sample (filled colored symbols with dotted guides to the eye) and Poisson distribution estimates (black; empty symbols are Monte-Carlo simulations, dashed lines are the SW analytic approximation of §C) with respect to their median are plotted as a function of the normalized radius $\tau\equiv\theta/\theta_{\rm 500}$ from the center of the cluster, against our nominal, normal-distribution approximation $S(\tau)$ (thin solid horizontal lines with left ticks). The anticipated VS region (the $2.52.25<\tau<$ 2.5 bin of previous $\gamma$-ray [reiss2018detection] and radio [HouEtAl23] stacked detections) is highlighted (vertical green shading).
3.4 Stacked source-excess models
For a significant $S>0$ excess, we use the TS-test to determine the CL at which the foreground-only null hypothesis is rejected in favor of one of two simple, effective models for the projected, dimensionless source-density $n(\tau)$: (i) a planar model, of a 2D ring around the cluster oriented within the plane of the sky; and (ii) a shell model, obtained by projecting a spherical 3D shell around the cluster.
For the planar model, we take $n(\tau)$ proportional to a 2D Gaussian of peak radius $\tau_{\rm sh}$ and width $\Delta\tau_{\rm sh}$, giving the three-parameter model
$n_{\rm{planar}}\left(\tau\right)\equiv\frac{n_{0}\tau_{\rm sh}}{\sqrt{8\pi\,\Delta\tau_{\rm sh}^{2}}}\exp\left[-\frac{\left(\tau-{\tau_{\rm sh}}\right)^{2}}{2\,\Delta\tau_{\rm sh}^{2}}\right]\mbox{ .}$ (3.24)
The prefactor is chosen such that in the $\Delta\tau_{\rm sh}\to 0$, thin-ring limit, a normalized solid-angle integration yields
[…]
sources inside the VS. Here, $d\omega$ is the solid-angle element $d\Omega$ normalized by either $\bar{\theta}_{500}^{2}$ (for SW) or $\theta_{500,{\mathsf{c}}}^{2}$ (CW), and $n_{0}$ is a constant dimensionless source density.
For the shell model, it suffices for our purposes to consider a zero-width 3D shell, giving upon projection the two-parameter model [e.g., keshet2018evidence]
$n_{\rm shell}\left(\tau\right)\equiv\frac{n_{0}\tau_{\rm sh}\Theta\left(\tau_{\rm sh}-\tau\right)}{2\sqrt{\tau_{\rm sh}^{2}-\tau^{2}}}\,,$ (3.26)
where $\Theta$ is the Heaviside step function. The normalization is chosen to again satisfy Eq. (3.25), to facilitate a meaningful comparison of $n_{0}$ in the two models.
3.5 Model fitting and TS-Test
We estimate the best-fit model parameters by a least-squares procedure. In general, one may minimize
$\chi^{2}\simeq\sum_{\tau,{\mathsf{c}}}\frac{\left[\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})-\mathcal{M}(\tau,{\mathsf{c}})\right]^{2}}{\mathcal{F}(\tau,{\mathsf{c}})+\mathcal{M}(\tau,{\mathsf{c}})}\mbox{ ,}$ (3.27)
where $\mathcal{M}(\tau,{\mathsf{c}})$ is the model excess at radial bin $\tau$ in cluster ${\mathsf{c}}$, and we exclude (henceforth) the innermost, $\tau<1.5$ bins, to avoid the noisy centers of clusters. However, clusters typically have some inner bins with no sources at all, so the above quantity is not well-approximated by a $\chi^{2}$ distribution. We therefore consider figures of merit based on the stacked quantities.
For SW stacking, we consider $\mathcal{N}(\tau)$, already co-added over clusters, as the relevant random variable, leading to
$\chi^{2}_{\rm\mbox{\tiny{SW}}}\simeq\sum_{\tau}\frac{\left[\mathcal{N}(\tau)-\mathcal{F}(\tau)-\mathcal{M}(\tau)\right]^{2}}{\mathcal{F}(\tau)+\mathcal{M}(\tau)}\mbox{ ,}$ (3.28)
where $\mathcal{M}\left(\tau\right)=\sum_{{\mathsf{c}}}\mathcal{M}(\tau,{\mathsf{c}})$ is accordingly the model excess co-added over clusters. For CW stacking, taking $S_{\rm CW}(\tau)$ as the relevant random variable yields
$\chi^{2}_{\rm\mbox{\tiny{CW}}}$ $\simeq\sum_{\tau}\frac{\left[\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})-\mathcal{M}(\tau,{\mathsf{c}})}{\sqrt{\mathcal{F}(\tau,{\mathsf{c}})}}\right]^{2}}{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\left[1+\frac{\mathcal{M}(\tau,{\mathsf{c}})}{\mathcal{F}(\tau,{\mathsf{c}})}\right]}\mbox{ .}$ (3.29)
In both methods, one may also sum over the two catalogs in order to fit both X-ray and radio signals simultaneously and estimate their joint significance.
Minimizing $\chi^{2}$ is equivalent in the normal-error approximation to finding the maximum $\mathcal{L}_{\rm max}$ of the likelihood $\mathcal{L}(\chi^{2})=\exp(-\chi^{2}/2)$. The significance of the excess within the model is estimated from the test statistic [MattoxEtAl96_TS]
$\mbox{TS}\equiv-2\ln\frac{\mathcal{L}_{\rm max,-}}{\mathcal{L}_{\rm max,+}}=\chi^{2}_{-}-\chi^{2}_{+}\,,$ (3.30)
where a subscript ‘--’ (‘$+$’) refers to the best-fitted model without (with) the modelled excess. Then TS has an approximate chi-squared distribution $\chi_{\mathsf{n}}^{2}$ of order $\mathsf{n}\equiv\mathsf{n}_{+}-\mathsf{n}_{-}$, i.e. with the number of degrees of freedom added by the excess-source model [Wilks1938]. Our single-parameter $1\sigma$ confidence intervals are defined through $\mbox{TS}=1$ after projecting onto the given parameter.
For modelling, we nominally adopt a $\Delta\tau=1/6$ resolution, to better capture the radial profile; results for other resolutions are given in §D. We use Monte-Carlo simulations to verify that although the data is not normally distributed, $\chi^{2}$ in both Eqs. (3.28) and (3.29) indeed follows the expected $\chi^{2}$ distribution (see §C), and so can be used to estimate the TS-based CLs, denoted ${\sigma_{\tiny{\mbox{TS}}}}$.
4 Virial excess and its implications
Our nominal stacking procedure yields the significance $S(\tau)$ profiles of the radially binned excess in Fig. 4, with the SW significance $S(\tau_{x},\tau_{y})$ of the 2D excess illustrated in Fig. 5. Fractional source overdensity $\delta(\tau)$ profiles are shown in Fig. 6, along with their effective modelling summarized in Table 2 and in more detail in Table 3. Each panel in these figures shows results for both X-ray and radio catalogs, as well as a joint analysis of both catalogs. Additional results and models are provided in §D.
4.1 Stacking
The stacking results are only weakly sensitive to the details of the analysis, as shown by the various sensitivity tests and method variations in §D, as well as by the similar results obtained in Figs. 4 and 6 using the SW (left panels) vs. CW (right panels) co-addition methods. In Fig. 5 and in the models of Table 2 and §4.2, we focus on the more conventional and well-calibrated SW stacking; CW models are deferred to Table 3.
As expected, the centers of the stacked clusters show a strong excess of both X-ray and radio sources associated with the ICM. This $\tau\lesssim 1.5$ excess reaches very high, $\delta\simeq 50$ ($\delta\simeq 20$) central overdensities in X-rays (radio), detected at extremely high, $S\simeq 100$ ($S\simeq 40$) CLs. These and all subsequent CLs include the aforementioned correction for Poisson-statistics.
We are more interested here in cluster peripheries, where a nearly identical, highly significant, $\delta\simeq 0.5$ overdensity is identified in both X-ray ($4.4\sigma$) and radio ($4.0\sigma$) sources. In both catalogs, this excess is highly localized, found in the same anticipated region (highlighted in Figs. 36), and peaked around the same $\tau\simeq 2.4$ normalized radius, identified previously as the VS in $\gamma$-ray [reiss2018detection] and radio [HouEtAl23] continuum stacking. Indeed, a joint-catalog analysis indicates a very high confidence, $5.9\sigma$ detection. Below, we discuss this signal, henceforth referred to as the VS excess, in detail.
In addition to the central and VS signals, both catalogs show evidence for a very peripheral, $5.5\lesssim\tau\lesssim 7.5$ excess. This weak, $\delta\simeq 0.2$ overdensity presents in both X-rays and radio at a low, $S\simeq 2$ significance in our nominal resolution scans. However, the signal is very extended, and so presents with a more significant, $\sim 2.5\sigma$ detection at low resolution. Furthermore, as the X-ray and radio signals overlap, a joint-catalog analysis implies a more significant, $\sim 3\sigma$ overdensity. Interestingly, a similar peripheral weak excess was previously identified independently, in continuum $\gamma$-ray, radio, and SZ analyses, suggesting non-spherical, elongated VS morphologies; see discussion in Ref. [reiss2018detection].
4.2 Modeling
Consider first the VS excess in X-ray sources. Fitting a planar (shell) model places the shock at $\tau_{\rm sh}=2.44^{+0.06}_{-0.19}$ ($2.50^{+0.00}_{-0.02}$), detected at a TS-significance of $3.8\sigma$ ($3.3\sigma$), as inferred from $\mbox{TS}\simeq 20.6$ with $\mathsf{n}=3$ degrees of freedom ($\mbox{TS}\simeq 14.2$ with $\mathsf{n}=2$). The width of the shock in the planar model is $\Delta\tau_{\rm sh}=0.03^{+0.04}_{-0.03}$; for a cluster of median sample parameters ($z\simeq 0.077$ and $\theta_{500}\simeq 0{\overset{{}^{\circ}}{.}}15$), this corresponds to only $\sim 20\mbox{ kpc}$. Given our prior knowledge of the VS radius, the effective significance of the detection is higher. For instance, if one adopts as priors a shock radius $\tau_{\rm sh}=2.4$ (based on [reiss2018detection] and [HouEtAl23]) and a sub-resolution shock width, then the TS-confidence rises to $4.5\sigma$ (for $\mathsf{n}=1$ remaining free parameter).
Modeling the radio-source excess yields very similar results. Here, the planar (shell) model indicates a shock of similar radii $\tau_{\rm sh}=2.44^{+0.04}_{-0.05}$ ($2.54^{+0.06}_{-0.03}$), detected at a $3.8\sigma$ ($3.3\sigma$) CL. The shock width in the planar model, $\Delta\tau_{\rm sh}=0.12^{+0.05}_{-0.04}$, is larger than inferred from X-rays, corresponding to $\sim 90\mbox{ kpc}$ for median (see Table 1) cluster parameters. Here too, if one adopts as priors $\tau_{\rm sh}=2.4$ and a sub-resolution shock width, the detection TS-confidence rises to $4.5\sigma$.
Figure 4: Significance $S(\tau)$ profiles of the SW/CW (left/right panel) co-added excess of X-ray (blue/cyan disks with solid guides to the eye), radio (red/orange squares with dashed guides), and co-added X-ray+radio (green triangles with dot-dashed guides) sources around our nominal MCXC cluster sample. Corrected confidence levels (thin, approximately horizontal curves) for X-ray (solid curves) and radio (dashed) are based on either a Poisson approximation (for SW, where they are accurate) or control clusters (for CW). The anticipated VS region is highlighted as in Fig. 3.
Figure 5: Excess-source significance $S$ in sky coordinates $\{\tau_{x},\tau_{y}\}$ for SW-stacked and folded (see labels) X-ray (folded onto left quadrant and octant), radio (bottom quadrant and octant) and joint (upper right quadrant) catalogs. Dotted circles at $\tau=2.5$, $4.5$, and $6.5$ are guides to the eye. Stacked source images around individual clusters are rotated randomly around the center of the cluster before cluster co-addition.
The X-ray and radio excess signals closely overlap, so a joint-catalog analysis produces similar results, at a higher CL. Here, the planar (shell) model indicates a $\tau_{\rm sh}=2.41^{+0.02}_{-0.02}$ ($\tau_{\rm sh}=2.50^{+0.82}_{-0.77}$) shock, detected at $5.3\sigma$ ($4.4\sigma$). The corresponding (planar model) shock width $\Delta\tau_{\rm sh}=0.10^{+0.03}_{-0.02}$ translates to $\sim 80\mbox{ kpc}$ for the median ($z\simeq 0.079$ and $\theta_{500}\simeq 0{\overset{{}^{\circ}}{.}}14$) parameters of the joint cluster sample. Adopting as priors a $\tau_{\rm sh}=2.4$ shock radius and a sub-resolution shock width would raise the TS-confidence to $5.8\sigma$.
Table 2: Nominal VS modeling.
Model Cat. $n_{0,X}^{*}$ $n_{0,\nu}^{*}$ $\tau_{\rm sh}$ $\Delta\tau_{\rm sh}$ TS ${\sigma_{\tiny{\mbox{TS}}}}$
(1) (2) (3) (4) (5) (6) (7) (8)
Shell 2RXS $8.2^{+1.4}_{-3.4}$ $2.50^{+0.00}_{-0.04}$ $14.2$ $3.3$
Shell NVSS $8.1^{+2.4}_{-2.3}$ $2.54^{+0.06}_{-0.03}$ $14.0$ $3.3$
Shell Joint $7.3^{+2.3}_{-2.1}$ $6.8^{+2.2}_{-2.1}$ $2.50^{+0.83}_{-0.78}$ $25.8$ $4.4$
Planar 2RXS $5.5^{+1.6}_{-1.5}$ $2.35^{+0.05}_{-0.02}$ $0.03^{+0.04}_{-0.03}$ $20.6$ $3.8$
Planar NVSS $6.5^{+2.1}_{-1.8}$ $2.44^{+0.04}_{-0.05}$ $0.12^{+0.05}_{-0.04}$ $20.4$ $3.8$
Planar Joint $5.8^{+1.7}_{-1.5}$ $5.9^{+1.9}_{-1.7}$ $2.41^{+0.02}_{-0.02}$ $0.10^{+0.03}_{-0.02}$ $38.1$ $5.3$
  • Nominal (SW; $\Delta\tau=1/6$) best-fitted parameters; additional models are provided in Table 3. Table columns: (1) Effective VS model (planar or shell); (2) Stacked catalog (2RXS, NVSS, or the two catalogs jointly); (3) and (4) Normalized, $n_{0}^{*}\equiv 10^{3}n_{0}$ X-ray and radio source densities in Eqs. (3.24) or (3.26); (5) Dimensionless shock radius; (6) Dimensionless signal width; (7) TS value; (8) TS-based significance.
Curiously, we find similar source densities for the VS excess in the X-ray and radio catalogs. With the planar model, for instance, $n_{0}\simeq 5\times 10^{-3}$ (sources per $\theta_{500}^{2}$) in X-rays and $n_{0}\simeq 7\times 10^{-3}$ in radio; the shell model yields $n_{0}$ values similar but slightly higher due to the 3D projection. The exact difference between X-ray and radio peak densities is not particularly meaningful, as $n_{0}$ is somewhat sensitive to details such as catalog parameters and cuts. Despite their similar properties, in §F we show that X-ray and radio excess signals are not dominated by the same individual objects.
Figure 6: Stacked fractional $\delta(\tau)$ overdensity (same notations as in Fig. 4), along with the best-fitted planar (thick curves, three-parameter) and shell (thin, two-parameter) models for the X-ray (solid curves) and radio (dashed) sources.
As argued in §1, an excess of AGN in a narrow radial range at the far outskirts of clusters would be highly unnatural. Our above results establish systematically and quantitatively that both X-ray and radio sources emerge in a very narrow, $\Delta r/r\lesssim 20$ radial range, at least in a large fraction of clusters. This sharp excess is inconsistent with enhanced AGN triggering, given the weak environmental gradients found at these large, $\gtrsim 2\mbox{ Mpc}$ radii, and the low ambient pressure even behind the VS. We conclude that the signals are unrelated to AGN; indeed, we find no significant localized signal when examining or stacking sources associated with identified AGN, as shown in §G. Moreover, the properties of the excess sources, discussed in §5, are inconsistent with AGN.
The robust spatial coincidence of the excess with the previously-determined VS location indicates that the excess is associated with the shock itself, and is most likely triggered by it. While the shock provides the strong spatial localization necessary to account for the sharpness of the excess, its downstream pressure is still far too weak to penetrate into the very center of a galaxy and trigger an AGN, so it is necessary to explore alternative models. Indeed, the constraints derived in §5 on the properties of the excess sources indicate that they constitute a different class of objects, explored in §6.
5 Properties of VS excess sources
The $\delta\simeq 0.5$ fractional overdensity of both X-ray and radio catalog sources within the narrow VS range indicates that roughly $\sim 1/3$ of the sources therein are associated with the VS, while the remaining $\sim 2/3$ majority are field sources. The sources found in the VS regions, listed in §H, are mostly unidentified; note that these peripheral sources lie well outside the radii of known ICM radio sources, including halos, relics, and even mega-halos [Keshet24], and cannot be attributed to the smooth distribution of cluster sources or to AGN (§G). While we cannot attribute individual sources to the VS, the most likely properties of those VS-induced sources can be deduced from a differential statistical analysis that compares the properties of sources in the VS region to their field counterparts.
Consider some source property ${\mathbb{P}}$ of interest, such as flux, spectral index, or size. Figures 714 compare, for various such properties, the distribution of ${\mathbb{P}}$ among sources in the VS region ($2.52.25<\tau<$ 2.5) to ${\mathbb{P}}$ among sources in the ICM region ($1.51.0<\tau<$ 1.5) and in the field. Here, we adopt a normalized, $357<\tau<$ 35 radial field range, instead of the fixed angular scale invoked earlier for field removal, in order to obtain a better behaved VS-to-field source-number ratio. In addition to the nominal analysis, we also consider different field regions ($3010<\tau<$ 30) and cuts that do not exclude the FB and Loop-I regions, in order to improve the statistics. For additional tests and source properties, see §E.
When relevant, we show analogous X-ray (left panels) and radio (right panels) ${\mathbb{P}}$ properties on the same figure. Each figure presents both differential $\mathbb{N}({\mathbb{P}})\equiv dN/d{\mathbb{P}}$ (top panels) and cumulative $N(<{\mathbb{P}})$ (bottom panels) distributions in the VS region (solid and dot-dashed green), compared to equivalent field (thick dashed and thin dotted red) and ICM (dot-dashed and double dot-dashed magenta) distributions scaled to the same $\Omega$. Bottom panels highlight also the cumulative, $N_{\mbox{\tiny VS}}(<{\mathbb{P}})-N_{\mbox{\tiny F}}(<{\mathbb{P}})$ distribution (solid blue) and the $S({\mathbb{P}})=[\mathbb{N}_{\mbox{\tiny VS}}-\mu(\mathbb{N}_{\mbox{\tiny F}})]/\sigma(\mathbb{N}_{\mbox{\tiny F}})$ local significance (dashed green) of the VS excess, with respect to the field distribution $\mathbb{N}_{\mbox{\tiny F}}$ inferred from $>200$ random field samples.
Figure 7 shows the logarithmic, ${\mathbb{P}}=\log_{10}F$ energy flux distribution. In X-rays, the VS excess arises mostly in the $2\times 10^{-13}\mbox{ erg}\mbox{ s}^{-1}\mbox{ cm}^{-2}\lesssim F_{X}\lesssim 2\times 10^{-12}\mbox{ erg}\mbox{ s}^{-1}\mbox{ cm}^{-2}$ flux range, dominating the cumulative distribution at local $\gtrsim 3\sigma$ significance levels, and appears unaffected by our higher $F_{X}$ upper limit. In radio, the excess arises mostly in the $10^{2}\mbox{ mJy}\lesssim F_{\nu}\lesssim 10^{3}\mbox{ mJy}$ range, with a local $\gtrsim 2\sigma$ significance increasing towards lower $F_{\nu}$, suggesting additional excess sources below our lower $F_{\nu}$ cutoff. The local excess of bright sources with $\sim 5$ Jy flux, seen in the figure, can probably be disregarded, as it is associated with only two sources and is not robust against changes in analysis details.
The cataloged sources do not have readily available associated redshifts. Assigning sources to the redshift of the nearby MCXC cluster, as seen in projection, produces an intrinsic, albeit noisy, estimate of the typical luminosity of excess VS sources. Figure 8 thus shows that the excess is associated with $10^{42}\mbox{ erg}\mbox{ s}^{-1}\lesssim L_{X}\lesssim 10^{43}\mbox{ erg}\mbox{ s}^{-1}$ 2RXS sources and with $10^{40}\mbox{ erg}\mbox{ s}^{-1}\lesssim\nu L_{\nu}\lesssim 10^{41}\mbox{ erg}\mbox{ s}^{-1}$ NVSS sources. (The luminosity of excess GMRT sources is a factor $\sim 2$ lower than in NVSS; see Fig. E.6.) Again, local excess signals at lower or higher luminosities cannot be substantiated, as they are associated with few sources and are not robust against changes to the analysis details.
Consider the (unlikely) possibility that the VS excess signals in X-rays and in radio arise from the same power-law spectrum, extending over more than eight orders of magnitude in photon energy (although the implied synchrotron origin of the X-rays is ruled out in §6). For the relevant VLA and ROSAT channels, the associated photon spectral index would then be
$\alpha\equiv-\frac{d\ln L}{d\ln\nu}\simeq 0.81\left(\frac{100\nu L_{\nu}}{L_{X}}\right)^{0.067}\,,$ (5.1)
providing a robust reference value for a spectral analysis.
Figure 7: Differential $\mathbb{N}({\mathbb{P}})$ (top panels) and cumulative $N(<{\mathbb{P}})$ (bottom panels with left axes) distributions of logarithmic, ${\mathbb{P}}=\log_{10}F$ energy flux in X-rays (left panels) and radio (right panels). Differential distributions are shown for the VS region (solid green) and comparable ICM (dot-dashed magenta) and field (red; thick dashed for mean value and thin dotted for $1\sigma$ dispersion) estimates; see text. Bottom panels show, in addition to the cumulative distributions in the VS region (dot-dashed green), field estimate (dotted red), and their difference (solid blue), also the local significance of the VS excess (dashed green with right axes). In radio, we add a $\sigma_{\rm smooth}^{2}\equiv(0.2\,F_{\nu})^{2}$ smoothing variance to the $F_{\nu}$ uncertainty, for visibility.
Figure 8: Distribution of luminosity, ${\mathbb{P}}=\log_{10}L_{X}$ in X-rays (left) and ${\mathbb{P}}=\log_{10}\nu L_{\nu}$ with added $\sigma_{\rm smooth}^{2}(L_{\nu})\equiv(0.2\,L_{\nu})^{2}$ variance (for visibility) in radio (right); notations are as in Fig. 7.
The spectral index $\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}$ is estimated by matching sources in the GMRT and NVSS catalogs through a projected proximity, $\theta<\theta_{\rm max}=30^{\prime\prime}$ angular separation criterion. Most ($\sim 90\%$) GMRT sources then show an NVSS counterpart, with only a small, $\pi\theta_{\rm max}^{2}N_{\rm NVSS}N_{\rm GMRT}/(4\pi f_{\Omega}N_{\rm pairs})\sim 1\%$ fraction of the resulting $N_{\rm pairs}\simeq 5.6\times 10^{5}$ pairs being false, chance associations. Here, $N_{\rm NVSS}\simeq 1.6\times 10^{6}$ and $N_{\rm GMRT}\simeq 5.8\times 10^{5}$ are respectively the numbers of NVSS and GMRT sources in the relevant, $f_{\Omega}=82\%$ fraction of the sky. Defining $\mathbb{P}=\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}$ for each pair, the radio VS excess may be inferred from Fig. 9 to arise mostly from $0.5\lesssim\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}\lesssim 0.8$ sources. The considerable $\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}<0.6$ excess suggests that this hard synchrotron power-law spectrum does not extend up to X-ray energies.
[…]
Figure 9: Distribution of $\mathbb{P}=\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}$ radio spectral indices among GMRT-NVSS catalog pairs; notations as in Fig. 7.
This estimate is broadly consistent but somewhat harder than in Eq. (5.1), and in the absence of NVSS-2RXS pairs (see §F), we cannot support such $\alpha_{\nu}$ values persisting out to X-rays. Moreover, while we cannot localize any excess in the X-ray spectral index $\Gamma_{X}$ obtained from power-law fits, both Mekal and black-body fits suggest $k_{B}T\gtrsim 1\mbox{ keV}$ excess-source temperatures, as shown in Figs. 10 and A.2, where $k_{B}$ is the Boltzmann constant. The combined implications of hard power-law spectra in radio and thermal spectra in X-rays are discussed in §6.
[…]
Figure 10: Distribution of the ${\mathbb{P}}=\log_{10}k_{B}T$ Mekal-fitted temperature; notations as in Fig. 7. Here we show also the cumulative ICM-source distribution (double-dot magenta).
Figure 11: Distribution of the ${\mathbb{P}}=\log_{10}|L|/I$ linearly-polarized radio fraction; notations as in Fig. 10.
This effect is quite robust, strengthening at high Galactic latitudes or at high declinations [where the typical Galactic Faraday rotation at NVSS frequencies drops below $\sim 30^{\circ}$; see, e.g., TaylorEtAl09], as illustrated by the figure. Such a polarization supports the association of the sources not only with the cluster, but also specifically with its VS, suggesting that the emission arises from a gaseous medium shocked by the VS, resulting in enhanced tangential magnetic fields. While the measured polarizations carry substantial uncertainties, the elevated polarized flux and the correlation between polarization vectors and radii among VS (but not field) sources suggest that the polarization effect is real.
Assigning again sources with nearby projected MCXC clusters, we may estimate the typical sizes of sources in the VS region. Figure 13 thus suggests that the X-ray excess is dominated by sources of $30\mbox{ kpc}\lesssim R\lesssim 100\mbox{ kpc}$ radii, whereas the radio excess arises from sources with semi-major axes in the range $10\mbox{ kpc}\lesssim a\lesssim 100\mbox{ kpc}$. An analogous examination of the semi-minor axes $b$ of the radio sources suggests that the VS excess arises from elongated sources, of $6\mbox{ kpc}\lesssim b\lesssim 30\mbox{ kpc}$. Similarly estimating the dimensionless (normalized to $R_{500}$) scales yields $0.05\lesssim\tau_{R}\lesssim 0.15$, $0.02\lesssim\tau_{a}\lesssim 0.2$, and $0.01\lesssim\tau_{b}\lesssim 0.15$, associated respectively with the X-ray radius and the radio semi-major and semi-minor axes (see §E). These scales are broadly consistent with the width $\Delta\tau_{\rm sh}$ of the stacked VS signal, as derived in §4.
The distribution of the angle $\varphi$ between the semi-major axis of a radio source and the line connecting it to the center of the cluster, presented in Fig. 14, appears bimodal. A significant excess is found for sources elongated both around $\varphi\simeq 30^{\circ}$, and perpendicular ($\varphi\simeq 90^{\circ}$) to the cluster radius; again, any well established $\varphi$ excess would support the association of the virial excess with the cluster, and specifically with the VS. The second, $\varphi\simeq 90^{\circ}$ excess (see also Fig. E.4) is associated with the more extended sources, which appear to be stretched along the VS; however, we cannot substantiate this excess, as it vanishes at high declinations as shown in the figure.
We examine the dependence of the VS excess upon cluster mass by crudely splitting the cluster sample into three $M_{500}$ bins, as shown in Fig. 15. The bins are adjusted to include a roughly fixed, $\sim 150$ number of clusters per bin, to provide comparable statistics in each bin; this is achieved only approximately, as we adopt the same mass bins for both X-ray and radio analyses. The figure suggests that the mass dependence of the VS excess is weak, unlike the strong mass dependence expected [LoebWaxman00, TotaniKitayama00, WaxmanLoeb00], simulated [KeshetEtAl03, KeshetEtAl04_EGRB], and indicated observationally [HouEtAl23] for smooth accretion.
[…]
Figure 12: Distribution of the $\mathbb{P}=|\phi|$ angle between polarization vector and cluster radius for all (left) and $\mbox{Dec}>20^{\circ}$ (right) sources; notations and sources as in Fig. 11. A $\sigma_{\rm smooth}^{2}(\varphi)=(5^{\circ})^{2}$ variance is added for visibility.
Figure 13: Distribution of source size: radius $\mathbb{P}=R$ in X-rays (left) and semi-major axis $\mathbb{P}=a$ with added $\sigma_{\rm smooth}^{2}(a)\equiv(0.2\,a)^{2}$ variance in radio (right) for visibility; notations are the same as in Fig. 7. Normalized (to $\theta_{500}$) and semi-minor axes are shown in Figs. E.2 and E.3.
Figure 14: Distribution of the $\mathbb{P}=|\varphi|$ angle between source elongation and cluster radius for all (left) and $\mbox{Dec}>20^{\circ}$ (right) sources; notations and sources as in Fig. 11. A $\sigma_{\rm smooth}^{2}(\varphi)=(5^{\circ})^{2}$ variance is added for visibility.
Figure 15: Significance $S(\tau)$ profile of excess X-ray (left panel) and radio (right) sources in different $M_{500}$ bins (see legend); other notations are the same as in Fig. 4.
6 Origin of the VS sources
The above results show a highly significant excess of X-ray and radio catalog sources narrowly peaked near the same, $\tau\simeq 2.4$ normalized radius where previously stacked $\gamma$-ray [reiss2018detection] and radio [HouEtAl23] continuum signals were attributed to VS-induced emission. This spatial coincidence indicates that the excess sources are directly associated with the VS of their host cluster, especially considering the very narrow, $\Delta\tau\simeq 0.1$ width of their stacked excess. Indeed, it is difficult to envision how a well-localized spike in source density at this particular radius could arise without involving the VS. The conclusion is further supported by anecdotal evidence, such as the well-defined angles of radio source polarization and possibly also elongation with respect to the host cluster. Here, we consider the physical mechanism by which the VS may induce the excess sources, as implied by the stacked signal (§4) and statistical source properties (§5).
The comparable and narrow widths of the X-ray and radio signals suggests that the sources radiate, or, more likely, brighten significantly, only during VS crossing or soon thereafter. For a body of length $l=100l_{100}\mbox{ kpc}$, such brightening would span a normalized $\Delta\tau\simeq 0.1l_{100}$ radial extent during the $t\simeq l/v_{s}\simeq 0.1l_{100}T_{1}^{-1/2}\mbox{ Gyr}$ passage of the body into the $k_{B}T_{d}\simeq T_{1}\mbox{ keV}$ ICM, approximating its infall through the VS as fixed at the $v_{s}\simeq(16k_{B}T/3\bar{m})^{1/2}$ shock velocity. This conclusion is supported by the $0.02\lesssim\{\tau_{R},\tau_{a}\}\lesssim 0.2$ extent of the individual sources comprising the excess, being comparable and not much smaller than the signal width.
The $\Delta P\lesssim 0.02\mbox{ eV}\mbox{ cm}^{-3}$ pressure change induced by the VS [e.g., ArnaudEtAl10, PlanckV13, HurierEtAl19, keshet20coincident] is much too weak to penetrate down to the vicinity of compact bodies, certainly those found in the centers of galaxies. As anticipated in §1, our results thus challenge previous claims that a well-localized excess of X-ray sources near the virial radius could be attributed to AGN. Indeed, sources already associated with identified AGN show no significant localized excess; see §G. Instead, sources triggered by the VS are likely to be extended and gaseous. Indeed, this conclusion is directly corroborated by the inferred, 10101010100100100100 kpc extent of the excess sources, as well as by the polarization angle of the radio emission.
[…]
$l_{\mbox{\tiny cool}}\simeq v_{s}t_{\mbox{\tiny cool}}\simeq v_{s}\frac{6\pi m_{e}cf_{B}}{\sigma_{T}\gamma_{\nu}B^{2}}\simeq 110\frac{\sqrt{B_{1}T_{1}}}{1+0.1B_{1}^{2}}\mbox{ kpc}$ (6.1)
is comparable to the width of the signal at the $B_{1}\equiv B/1\mbox{ $\mu$G}\simeq 1.9$ maximum of the cooling time $t_{\mbox{\tiny cool}}(B)$. Here, $c$ is the speed of light, $\sigma_{T}$ is the Thomson cross-section, $m_{e}$ is the electron mass, and we defined the fractions $f_{\epsilon}\equiv 1-f_{B}$ and
$f_{B}\equiv\frac{B^{2}/8\pi}{u_{cmb}+B^{2}/8\pi}\simeq\frac{0.1B_{1}^{2}}{1+0.1B_{1}^{2}}$ (6.2)
of IC and synchrotron CRE emission, $u_{cmb}$ being the CMB energy density.
[…]
In contrast to radio, the $\epsilon=\epsilon_{1}\mbox{ keV}$ X-ray sources are unlikely to arise from non-thermal, neither IC nor synchrotron, emission. Indeed, IC emission over the relevant scales, dominated by $\gamma_{\epsilon}\simeq 10^{3}\epsilon_{1}^{1/2}$ CREs scattering CMB photons, would extend over a far too long, $l_{\mbox{\tiny cool}}\simeq 2(T_{1}/\epsilon_{1})^{1/2}\mbox{ Mpc}$ cooling length, greatly exceeding the width of the signal; IC emission is also too faint to explain the sources, as shown below. The opposite problem of $l_{\mbox{\tiny cool}}\lesssim 9(T_{1}B_{1}/\epsilon_{1})^{1/2}\mbox{ pc}$ cooling and $l_{\mbox{\tiny acc}}<cl_{\mbox{\tiny cool}}/v_{s}$ acceleration lengths much too short to explain the signal would arise if X-rays were synchrotron emission, also requiring a radio spectrum similar to Eq. (5.1), which is softer than inferred in §5, as well as an ongoing acceleration of excessively energetic, $\gamma\gtrsim 10^{8.5}(\epsilon_{1}/B_{1})^{1/2}$ CREs. In either case, if CREs were responsible for the X-ray sources, we would expect to find at least a few 2RXS–NVSS pairs, absent in the present data.
Therefore, the X-ray excess is likely dominated by thermal emission from hot plasma. Indeed, the X-ray spectral fits are marginally more consistent with excess sources having high, $k_{B}T\gtrsim 1\mbox{ keV}$ temperatures, than with sources of power-law spectra (see Figs. 10, A.1, and A.2). In addition, the evidence suggests (albeit at low confidence) that the radio signal has a radial extent $\sim 2.5$ times wider than in X-rays (see Table 2) and that the radio $a$ and $b$ distributions extend to scales somewhat larger than the X-ray radius $R$ (see Figs. 13, E.2, and E.3), consistent with a thermal X-ray component being more compact than its relativistic, radio counterpart.
[…]
Such a shock, heating a fraction $f_{s}$ of the gas mass $M_{g}$ and depositing a fraction $\xi_{e}$ of the resulting thermal energy in CREs, leads to a synchrotron luminosity
$\nu L_{\nu}$ $\simeq$ $\frac{2\alpha-1}{2}\gamma_{\nu}^{-(2\alpha-1)}\xi_{e}f_{B}\frac{v_{s}}{\Delta r(\nu)}\frac{f_{s}M_{g}}{m_{p}}k_{B}T$
$\simeq$ $10^{40}\xi_{1}f_{B}\left(\frac{f_{s}M_{g}}{10^{11}M_{\odot}}\right)\frac{T_{1}^{3/2}B_{1}^{0.05}}{\Delta r_{100}}\mbox{ erg}\mbox{ s}^{-1}\,,$
where we defined $\xi_{1}\equiv\xi_{e}/1\%\simeq 1$ based on detected continuum VS signals [KeshetEtAl17, reiss2018detection, keshet2018evidence, keshet20coincident, HouEtAl23], and $\Delta r_{100}\equiv\Delta r/100\mbox{ kpc}\simeq 1$ based on our NVSS stacking. Equation (6) can explain at least the fainter of the $\nu L_{\nu}\simeq 10^{40\mbox{\scriptsize{--}}41}\mbox{ erg}\mbox{ s}^{-1}$ sources, even without accounting for a compressed or re-accelerated pre-existing CRE population. As mentioned above, the counterpart IC signal (replacing $f_{B}\to f_{\epsilon}$ in Eq. 6) is much too faint to explain $L_{X}$ for plausible, $B\gtrsim 0.1\mbox{ $\mu$G}$ fields.
An interesting possibility is that the excess sources, or a subset of these sources, are associated with remnant lobes produced by galactic outflows from previous AGN or star formation bursts. Such lobes could provide seed CREs and magnetic fields, more easily explaining the excess radio sources even at their high-luminosity end. For example, a VS-induced, factor $\sim 10$ brightening of faint lobes could account for the entire radio excess, if they comprise $\sim 10\%$ of the rapidly declining (without the $F_{\nu}^{*}$ cut) $\mathbb{N}(L_{\nu})$ distribution of field sources. Remnant galactic outflows could better explain why in the radio, the radial width of the stacked excess and the high end of the source-size distribution are somewhat larger than in X-rays, and may possibly account for the elongation angle $\varphi$ distribution in Fig. 14 if the lobes are, e.g., longer than $l_{\mbox{\tiny cool}}$ or are stronger along a large-scale filament feeding the cluster. While the X-ray sources should still have a thermal origin, they could be strengthened if such outflows enrich their host halos.
7 Summary and discussion
Following reports of excess X-ray sources in galaxy-cluster peripheries, we stack X-ray (2RXS) and radio (NVSS, GMRT, VLSSr) catalog sources around MCXC galaxy clusters. We impose cuts on the MCXC to match source-catalog limitations (Table 1), and an upper (lower specific) flux cutoff on the X-ray (radio) catalog, to avoid the dominant bright (faint) sources in the foreground (background); see Fig. 2. Our results are not, however, sensitive to any of our cuts or analysis details (see §D).
The radially binned and stacked projected source-density profiles are very similar to each other in X-rays and in radio (see Fig. 6, the fit results in Table 2, and their counterparts in §D), presenting a pronounced peak near the same, $\tau\simeq 2.4$ normalized radius where source-masked $\gamma$-ray [reiss2018detection] and radio [HouEtAl23] continuum signals were previously detected and attributed to emission from the VS. Both X-ray and radio catalog signals are highly localized ($\Delta\tau\lesssim 0.1$), indicating that the excess sources are directly driven by the VS.
These excess X-ray ($\sim 4\sigma$ CL for $443$ clusters), radio ($\sim 4\sigma$ for $485$ clusters), and joint ($>5\sigma$) source signals are significant enough (even without priors; see §4 and §D) to see directly in the stacked images even without full radial binning (Figs. 5 and D.7). The signals present robustly, independently, and with similar parameters in the four (2RXS, NVSS, GMRT, and VLSSr) catalogs, and are not sensitive to variations in our analysis methods and parameters (see §D). The signals and their CLs are verified using control cluster samples, showing that the stacked excess closely follows the expected Poisson statistics (Fig. 3 and §C) away from the real clusters.
The properties of the excess VS sources can be deduced only statistically and separately in each catalog, because the individual sources comprising the excess cannot be isolated from the $\sim$ twice as large number of coincident field sources, and we cannot demonstrate a VS excess of X-ray–radio source pairs (see §F) with present statistics. We thus characterize the excess distributions of the flux (Fig. 7), luminosity (Fig. 8), and size (Fig. 13) of the sources, the temperature (Fig. 10) of the X-ray sources, and the spectral index (Fig. 9) and elongation angle (Fig. 14) of the radio sources, as well as their linear polarization fraction (Fig. 11) and angle (Fig. 12). Additional source properties are provided in §E.
The inferred properties of the excess X-ray and radio sources support their association with the VS, and constrain the physical mechanism responsible for their emergence (see §6). In particular, the radial, $\phi\simeq 0$ polarization of the radio sources, and their elongation angle distribution with its tentative tangential, $\varphi\simeq 90^{\circ}$ component, relate these sources to the VS orientation. The hard, $0.55\lesssim\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}\lesssim 0.8$, partially polarized radio signal is identified as synchrotron emission from CREs energized by the VS, gyrating in predominately tangential magnetic fields amplified by the VS, whereas the X-ray emission is most probably thermal, arising from an extended plasma distribution, compressed and heated to $k_{B}T\gtrsim 1\mbox{ keV}$ temperatures by the VS. The excess sources are thus identified as heated gas clumps accreted through the VS, probably the hot halos of galaxies or galaxy aggregates, possibly enriched by dormant lobes from galactic outflows.
The narrowness of the excess-source signals and their coincidence with each other and with the VS also corroborate previous claimed detections of VS signals based on continuum emission [KeshetEtAl17, reiss2018detection, keshet2018evidence, keshet20coincident, HurierEtAl19, HouEtAl23], which were more complicated due to smaller numbers of relevant clusters, as well as challenges associated for example with data reduction, point-source removal, and interferometric sidelobes. A comparison with these studies indicates that in all cases, planar models fit the stacked $\tau\sim 2.4$ signals better than their projected-shell counterparts, and that a secondary weaker, broader, and farther, $5\lesssim\tau\lesssim 7$ signal emerges (cf. Ref. [reiss2018detection], [HouEtAl23], and the present Figs. 46). Both effects indicate that a spherical, isolated model of the VS is oversimplified. The secondary signal would imply, if established at high confidence, that projected VS are elongated, with a mean $\sim 2.5$ aspect ratio, as indicated in Coma [KeshetEtAl17, keshet2018evidence], while the better fit of the planar model may also reflect quenching of the background by the VS; for a discussion, see Ref. [HouEtAl23]. For the present signals, comprised of discrete excess sources, a planar model may also be favored by catalog selection effects preferring VS-compressed sources seen edge-on, but we are unable to study this possibility or the nature of the $\tau\sim 6$ excess sources with the available statistics.
Better statistics and new constraints are expected with future source and cluster catalogs, and are already accessible in part with catalogs recently made available. Following this work, an analysis [IlaniEtAl24] of the eROSITA-DE Early Data Release (EDR) catalogs [BrunnerEtAl22eRosita, LiuEtAl22eRosita, BaharEtAl22eRosita] established a significant, $3\sigma$$4\sigma$ excess of X-ray sources at $2.252.0<\tau<$ 2.25 radii around EDR clusters. In radio, the deep LOFAR Two-metre Sky Survey [LoTSS; Shimwell19] and Rapid ASKAP Continuum Survey [RACS; McConnellEtAl20] can be used to test our results, supplement them with better statistics, stronger spectral constraints, and additional polarization information, and possibly extend them in the future to higher redshifts.
A study of individual VS-driven source candidates is beyond the scope of the present work, but should become feasible with future dedicated studies and upcoming catalogs. In particular, excess VS-driven radio sources, which should become clearer and more radially elongated at very low frequencies due to the longer $l_{\mbox{\tiny cool}}\propto\nu^{-1/2}$, may have already been observed, for example as the “accretion relic” or the NGC 4849 signature in Coma [BonafedeEtAl22], both extending out to the IC-traced elliptical VS [KeshetEtAl17, keshet2018evidence]. In the absence of concrete examples of VS-driven sources or an excess of X-ray–radio pairs, our statistical determination of the excess-source properties relies on a fairly small number of sources, and may be confused by superimposing different types of sources. Thus, the physical origins of the X-ray and radio sources may differ more significantly from each other than inferred above, and tentative bimodality signatures (e.g., in $a$, $b$, $\varphi$, and $T$) may indicate multiple source types within each catalog.
The excess radio sources should emit an IC signal at hard X-ray and $\gamma$-ray energies $\epsilon$. In the strong, $p\to 2$ shock limit, the anticipated IC luminosity of such a source is $\epsilon L_{\epsilon}^{\mbox{\tiny(IC)}}\simeq(f_{\epsilon}/f_{B})\nu L_{\nu}\simeq 10^{40\mbox{\scriptsize--}41}B_{1}^{-2}\mbox{ erg}\mbox{ s}^{-1}$ (assuming $l_{\mbox{\tiny cool}}\simeq\Delta r$ for NVSS sources, as inferred in §6). Such values are comparable, for $B\simeq 1\mbox{ $\mu$G}$ fields within the source, to the luminosity of the entire VS ring, as inferred from $\gamma$-rays in stacked [reiss2018detection] and individual [KeshetEtAl17, keshet2018evidence, keshet20coincident] clusters. Hence, a few such sources, activated by some VS, may outshine the entire IC signal from smooth accretion through this VS, especially at the lower energies $\epsilon$ accessible by the $p>2$ source CREs. Similarly, a $\nu L_{\nu}\simeq 10^{40\mbox{\scriptsize{--}}41}\mbox{ erg}\mbox{ s}^{-1}$ radio source outshines the $\nu L_{\nu}\simeq 10^{39}\mbox{ erg}\mbox{ s}^{-1}$ synchrotron ring from smooth accretion (extrapolating the Ref. [HouEtAl23] signal to higher frequencies with their median parameters).
Nevertheless, the radiative VS signatures due to excess sources vs. smooth accretion are easily distinguished from each other by their markedly different properties, in particular brightness, extent, spectra, and cluster-mass dependence. For instance, in contrast to the strong $M_{500}$ dependence anticipated and measured [HouEtAl23] for smooth accretion, we find no significant mass dependence in the excess-source properties, as demonstrated in Fig. 15. Moreover, localized sources were vigorously removed and masked in previous Fermi-LAT and LWA analyses. In contrast, a discrete source component cannot be ruled out in the somewhat high [keshet20coincident] VS flux inferred in Coma from VERITAS data [KeshetEtAl17].
The link established here between the VS and a coincident spike in discrete thermal (X-ray) and non-thermal (radio, and possibly also X-ray and $\gamma$-ray) sources opens new pathways to tracing the VS, mapping its relation to the cosmic web, and uncovering the underlying physical processes. If the multiple, bright peripheral X-ray sources per cluster reported previously at $z\sim 1$ are indeed identified with the VS, then the redshift evolution of VS sources is substantial, and similar analyses could shed light on the cosmological evolution of clusters, galaxies, galactic outflows, or other infalling objects. The strong, naively $L_{X}\propto\rho^{2}\propto(1+z)^{6}$ redshift dependence of thermal emission from shocked objects would render them useful VS tracers at high redshifts. If the weak dependence of the signal upon host cluster, hinted by Fig. 15, is confirmed, then a radial spike in discrete sources around an object could map its VS even on small, sub-cluster scales.
Acknowledgements
We thank T. Boller, I. Rofeh, D. Kushnir, I. Gurwich, I. Reiss, Y. Lyubarsky, M. Gedalin, and Y. Faerman, for helpful discussions and support. This research has received funding from the Israel Science Foundation (ISF grants No. 1769/15 and 2126/22), from the IAEC-UPBC joint research foundation (grant No. 300/18), and from the Ministry of Science, Technology & Space, Israel.
May Gideon Ilani’s memory be a blessing.
Supplementary Material
Appendix A Spectral modeling
Figure A.1 shows that while the radio spectral index has a significant $\sim 5\sigma$ excess of $\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}\simeq 0.55$ indices associated with $\sim 10$ sources, and, in addition, indications of $\sim 10$ more $0.6\lesssim\alpha_{150\mbox{ \scriptsize MHz}}^{1.4\mbox{ \scriptsize GHz}}\lesssim 0.8$ excess sources (right panel), there is no consistent signal in the $\Gamma_{X}$ distribution of X-ray spectral indices in the power-law fit (left panel). The temperatures of both Mekal and black-body X-ray spectral fits in Fig. A.2 show a more localized excess, consistent with about half of the excess sources having $T\gtrsim 1\mbox{ keV}$ temperatures.
Figure A.1: Distributions of power-law spectral indices ${\mathbb{P}}=\Gamma_{X}$ in X-rays (left panel) and $\mathbb{P}=\alpha_{150\mbox{\scriptsize MHz}}^{1.4\mbox{\scriptsize GHz}}$ in radio (right); notations are the same as in Fig. 7.
Figure A.2: Distribution of X-ray Mekal (left) and black-body (right) fitted temperature $\mathbb{P}=\log_{10}T$. Notations are as in Fig. 11.
Appendix B Catalog source sizes
The 2RXS catalog provides an estimate of the angular extent of each source, expressed (T. Boller, private communications 2020) as the radius of the source minus the radius of the point spread function (PSF). For simplicity, we approximate the source radius as the e-fold drop in photon distribution [boese2001maximum]. The PSF varies with the location of the source across the PSPC field of view, so we adopt an averaged, approximate PSF value. For this purpose, we take the e-fold radius of the RASS PSF model [boese2000rosat], providing an estimated radius $r_{psf}\sim 40^{\prime\prime}$. We verify that this result is sensible, by examining individual sources. Our conclusions are not sensitive to the precise value of the actual PSF.
[…]
Appendix C Confidence levels for Poisson statistics
Monte-Carlo simulations were used to numerically estimate the CLs $S_{p}$ for Poisson, rather than normal, statistics, demonstrated in Fig. 3 (the empty black symbols show the $S$ values corresponding to $S_{p}=\pm 1,\pm 2\ldots$), but it is useful to derive an approximate expression for $S$ given $S_{p}$. For a Poisson distribution of mean $\lambda$, where measurement $k$ has probability $p(k)=e^{-\lambda}\lambda^{k}/k!$, we may analytically sum the probabilities of equal or larger (smaller) $k$ values, in order to quantify the significance of an excess (deficit) in units of Gaussian standard errors,
$\!\!\!\!\!\!S_{p}(k;\lambda)=\begin{cases}\vspace{0.1cm}\sqrt{2}\,\mathrm{erfc}^{-1}\left[\frac{\Gamma(k)-\Gamma(k,\lambda)}{\Gamma(k)/2}\right]>0&\mbox{if }k>\lambda\,;\\ \vspace{0.1cm}-\sqrt{2}\,\mathrm{erfc}^{-1}\left[\frac{\Gamma(1+k,\lambda)}{\Gamma(1+k)/2}\right]<0&\mbox{if }k<\lambda\,;\\ 0&\mbox{if }k=\lambda\,,\end{cases}$ (C.1)
where $\mathrm{erfc}(x)$ and $\Gamma(x)$ are respectively the complementary error and Euler Gamma functions. For SW co-addition, where $S_{\rm\mbox{\tiny{SW}}}$ is the function of $\mathcal{N}$ and $\mathcal{F}$ given in Eq. (3.13), one can thus relate $S_{\rm\mbox{\tiny{SW}}}$ to $S_{p}$ for any given $\mathcal{F}$ by solving the equation
$S_{p}=S_{p}(\mathcal{N}=\mathcal{F}+S_{\rm\mbox{\tiny{SW}}}\sqrt{\mathcal{F}};\mathcal{F})\mbox{ .}$ (C.2)
The resulting $S_{\rm\mbox{\tiny{SW}}}$ profiles are shown for integer $S_{p}$ values in the SW figures (as dashed black curves).
For modelling, we Monte-Carlo simulate our $\chi^{2}_{\rm\mbox{\tiny{SW}}}$ and $\chi^{2}_{\rm\mbox{\tiny{CW}}}$ expressions under the field-only null hypothesis and Poisson statistics. We verify that the results follow the expected $\chi^{2}$ distributions, thus justifying the parameter fit and TS significance estimates. In particular, when fitting models to the stacked excess, we estimate the relative goodness of fit using the TS test, following Eqs. (3.28)–(3.30) and Wilks’ theorem [Wilks1938], which are based on underlying normal statistics. To verify that the results do not change significantly under Poisson statistics, we use Monte-Carlo simulations, with the null hypothesis of Poisson-distributed field sources only, to estimate $\chi^{2}$ and TS. The resulting distributions are found to closely follow the anticipated $\chi_{\mathsf{n}}^{2}$ distribution with the correct effective number $\mathsf{n}$ of degrees of freedom.
Appendix D Convergence and sensitivity tests
Our results are qualitatively insensitive to variations in the selection criteria (of clusters, sources, or field sources) and analysis methods (of field removal and source binning, stacking, or modeling). This robustness is demonstrated in Figs. D.1D.7 below, each varying some analysis parameter with little qualitative impact on the outcome with respect to the nominal results (which are shown for reference as a solid red curve in most figures).
In particular, the results are shown to be robust to variations in the $|b|$ cut (Fig. D.1), the masking of clusters in different sky regions and the minimal projected cluster interdistance (Fig. D.2), the order of the polynomial fit to the field source distribution $\mathcal{F}$ around each cluster (Fig. D.3), the radial $\Delta\tau$ resolution (Fig. D.4), the X-ray and radio flux cutoffs $F^{*}$ (Fig. D.5), the likelihood $\mathcal{L}$ cutoff on the X-ray sources (Fig. D.6), and the presentation and folding of stacked sources in 2D sky coordinates (e.g., varying the resolution and smoothing of the map; Fig. D.7). When adopting the nominal, constant density $\mathcal{F}$ of field sources (i.e. a polynomial of order zero) around each cluster, we vary the region used to measure $\mathcal{F}$, with inner radii spanning the range $1^{\circ}{\mbox{--}}5^{\circ}$ and outer radii in the range $6^{\circ}{\mbox{--}}10^{\circ}$, resulting in only negligible changes to the $S(\tau)$ profile.
The same procedure used to test for a mass dependence in Fig. 15 is used in Fig. D.8 to examine the dependence of the excess upon $\theta_{500}$, here with six bins of $\sim 150$ clusters each, extending beyond the nominal $\theta_{500}$ range. For sufficiently extended clusters, the $\theta_{500}$ dependence appears to be weak. However, the VS excess entirely vanishes for compact, $\theta_{500}\lesssim 0{\overset{{}^{\circ}}{.}}1$ clusters, at least in X-rays. The high resolutions of the catalogs suggest that this disappearance is association with the intrinsic widths of the signals. Indeed, this cutoff scale corresponds to $d_{A}(\bar{z})\theta_{500}\Delta\tau\simeq 140\mbox{ kpc}$, broadly consistent with the signal widths inferred in §4. Here, $d_{A}$ is the angular diameter distance, and we adopted the median $\bar{z}\simeq 0.08$ redshift in the sample.
Variations in the modelling of the VS excess are demonstrated in Table 3, expanding on the nominal models presented in Table 2. Here, we also examine the use of priors on the model (numbers in bold are frozen parameters). As the table shows, the determination of the position and width of the VS excess is robust, and the significance of the detection increases if prior information, in particular regarding the shock position, is used.
Figure D.1: Significance $S(\tau)$ profiles of the SW-stacked X-ray (left panel) and radio (right panel) source excess under variations in the Galactic latitude cutoff $|b|$ (see legend); the nominal analysis is shown, for reference, as a solid red curve. The anticipated VS region is highlighted as in Fig. 3. Notice the contribution of $|b|<10^{\circ}$ radio sources, neglected in the nominal analysis.
[…]
Figure D.2: Variations in cluster masking in different regions (see legend), with neighboring clusters defined as pairs within $5{\theta_{500}}$ in projection; notations are the same as in Fig. D.1.
Figure D.3: Same as the nominal Fig. 4, but with field source densities $\mathcal{F}$ inferred from a fourth-order polynomial fit to their sky distribution within $1\leq\tau\leq 15$ around each cluster.
Figure D.4: Same as the nominal Fig. 4, comparing $\Delta\tau=1/4$ (top panels) and $\Delta\tau=1/6$ (bottom) resolutions.
Figure D.5: Variations in flux cutoffs (see legend), including the case of no flux cut at all (orange diamonds); notations are the same as in Fig. D.1.
Figure D.6: Variations in X-ray source likelihood (see legend); notations are the same as in the left panel of Fig. D.1.
Figure D.7: Same as Fig. 5, but for $\Delta\tau=1/2$ resolution (left panel) or after convolution with a Gaussian matrix of $0.5\theta_{500}$ standard deviation (right panel).
Figure D.8: Same as Fig. 15, but for different ${\theta_{500}}$ bins (see legend).
Table 3: Best-fitted parameters for different variants of the virial-excess model.
Method $\Delta\tau$ Catalog Model $\mathsf{n}$ $n_{0,x}$ $n_{0,r}$ $\tau_{sh}$ $\Delta\tau_{sh}$ TS ${\sigma_{\tiny{\mbox{TS}}}}$
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11)
SW $1/4$ 2RXS Shell 2 $7.0^{+2.2}_{-2.0}$ $2.50^{+0.00}_{-0.02}$ $14.0$ $3.32$
Planar 1 $5.4^{+1.5}_{1.5}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $22.8$ $4.77$
2 $5.4^{+1.6}_{-1.1}$ $\mathbf{2.40}$ $0.05^{+0.04}_{-0.05}$ $22.8$ $4.39$
3 $5.4^{+2.1}_{-1.7}$ $2.44^{+0.06}_{-0.19}$ $0.04^{+0.05}_{-0.04}$ $22.8$ $4.08$
NVSS Shell 2 $8.4^{+2.5}_{-2.4}$ $2.52^{+0.05}_{-0.02}$ $14.7$ $3.42$
Planar 1 $6.9^{+1.9}_{-1.7}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $21.3$ $4.61$
2 $6.9^{+1.4}_{-1.4}$ $\mathbf{2.40}$ $0.12^{+0.04}_{-0.04}$ $21.3$ $4.23$
3 $6.6^{+2.1}_{-2.0}$ $2.44^{+0.06}_{-0.06}$ $0.10^{+0.05}_{-0.10}$ $21.7$ $3.96$
Joint Shell 3 $7.0^{+2.2}_{-2.0}$ $6.9^{+0.7}_{-0.7}$ $2.50^{+0.24}_{-0.25}$ $27.8$ $4.61$
Planar 2 $6.1^{+1.2}_{-1.2}$ $5.9^{+1.2}_{-1.2}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $42.6$ $6.20$
3 $6.1^{+1.2}_{-1.2}$ $6.9^{+0.7}_{-0.7}$ $\mathbf{2.40}$ $0.09^{+0.03}_{-0.03}$ $42.6$ $5.93$
4 $5.9^{+1.7}_{-1.5}$ $5.7^{+1.7}_{-1.6}$ $2.47^{+0.02}_{-0.02}$ $0.04^{+0.02}_{-0.02}$ $43.3$ $5.75$
$1/6$ 2RXS Shell 2 $8.2^{+1.4}_{-3.4}$ $2.50^{+0.00}_{-0.04}$ $14.2$ $3.34$
Planar 1 $5.3^{+1.5}_{-1.5}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $19.8$ $4.45$
2 $5.3^{+1.6}_{-1.6}$ $\mathbf{2.40}$ $0.07^{+0.03}_{-0.03}$ $19.8$ $4.06$
3 $5.5^{+1.6}_{-1.5}$ $2.35^{+0.05}_{-0.02}$ $0.03^{+0.04}_{-0.03}$ $20.6$ $3.83$
NVSS Shell 2 $8.1^{+2.4}_{-2.3}$ $2.54^{+0.06}_{-0.03}$ $14.0$ $3.32$
Planar 1 $6.9^{+1.9}_{-1.8}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $19.8$ $4.45$
2 $6.9^{+2.1}_{-1.4}$ $\mathbf{2.40}$ $0.14^{+0.04}_{-0.04}$ $19.8$ $4.06$
3 $6.5^{+2.1}_{-1.8}$ $2.44^{+0.04}_{-0.05}$ $0.12^{+0.05}_{-0.04}$ $20.4$ $3.81$
Joint Shell 3 $7.3^{+2.3}_{-2.1}$ $6.8^{+2.2}_{-2.1}$ $2.50^{+0.83}_{-0.78}$ $25.8$ $4.41$
Planar 2 $5.8^{+1.2}_{-1.2}$ $5.9^{+1.2}_{-1.2}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $38.0$ $5.83$
3 $5.8^{+1.2}_{-1.2}$ $5.9^{+1.8}_{-1.2}$ $0.10^{+0.03}_{-0.02}$ $\mathbf{2.40}$ $38.0$ $5.55$
4 $5.8^{+1.7}_{-1.5}$ $5.9^{+1.9}_{-1.7}$ $2.41^{+0.02}_{-0.02}$ $0.10^{+0.03}_{-0.02}$ $38.1$ $5.31$
CW $1/4$ 2RXS Shell 2 $4.2^{+2.1}_{-1.9}$ $2.50^{+0.02}_{-0.05}$ $5.9$ $1.94$
Planar 1 $3.7^{+1.4}_{-1.1}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $12.1$ $3.48$
2 $3.7^{+1.5}_{-1.1}$ $\mathbf{2.40}$ $0.01^{+0.03}_{-0.01}$ $12.1$ $3.04$
3 $4.3^{+1.7}_{-1.5}$ $2.47^{+0.03}_{-0.21}$ $0.04^{+0.05}_{-0.04}$ $12.8$ $2.80$
NVSS Shell 2 $7.6^{+2.4}_{-2.2}$ $2.50^{+0.03}_{-0.03}$ $18.9$ $3.95$
Planar 1 $6.3^{+1.7}_{-1.5}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $24.9$ $4.99$
2 $6.3^{+1.9}_{-1.3}$ $\mathbf{2.40}$ $0.11^{+0.03}_{-0.03}$ $24.9$ $4.62$
3 $6.2^{+1.9}_{-2.0}$ $2.41^{+0.08}_{-0.06}$ $0.11^{+0.04}_{-0.10}$ $24.9$ $4.32$
Joint Shell 3 $4.9^{+2.4}_{-2.1}$ $7.7^{+0.7}_{-0.8}$ $2.50^{+0.24}_{-0.25}$ $27.5$ $4.58$
Planar 2 $3.8^{+1.2}_{-1.2}$ $5.9^{+1.8}_{-1.2}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $36.9$ $5.74$
3 $3.8^{+1.2}_{-1.2}$ $7.7^{+0.7}_{-0.8}$ $\mathbf{2.40}$ $0.10^{+0.03}_{-0.03}$ $36.9$ $5.46$
4 $3.8^{+1.4}_{-1.2}$ $5.5^{+2.1}_{-1.6}$ $2.47^{+0.02}_{-0.02}$ $0.04^{+0.02}_{-0.02}$ $37.5$ $5.26$
$1/6$ 2RXS Shell 2 $4.3^{+2.0}_{-1.8}$ $2.50^{+0.01}_{-0.04}$ $6.7$ $2.10$
Planar 1 $3.8^{+1.5}_{-1.2}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $12.0$ $3.47$
2 $3.8^{+1.5}_{-1.2}$ $\mathbf{2.40}$ $0.07^{+0.04}_{-0.07}$ $12.0$ $3.03$
3 $3.9^{+1.7}_{-1.5}$ $2.39^{+0.10}_{-0.06}$ $0.07^{+0.04}_{-0.07}$ $12.1$ $2.69$
NVSS Shell 2 $7.6^{+2.3}_{-2.2}$ $2.51^{+0.03}_{-0.01}$ $16.7$ $3.68$
Planar 1 $6.3^{+1.7}_{-1.6}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $22.0$ $4.69$
2 $6.3^{+1.9}_{-1.3}$ $\mathbf{2.40}$ $0.14^{+0.05}_{-0.04}$ $22.0$ $4.30$
3 $6.2^{+2.1}_{-1.8}$ $2.42^{+0.05}_{-0.05}$ $0.13^{+0.05}_{-0.04}$ $22.1$ $4.00$
Joint Shell 3 $4.8^{+2.4}_{-2.1}$ $7.2^{+0.7}_{-0.7}$ $2.51^{+0.24}_{-0.25}$ $25.0$ $4.33$
Planar 2 $3.8^{+1.1}_{-1.1}$ $5.6^{+1.1}_{-1.1}$ $\mathbf{2.40}$ $\mathbf{0.01}$ $33.2$ $5.42$
3 $3.8^{+1.1}_{-1.1}$ $7.2^{+0.7}_{-0.7}$ $\mathbf{2.40}$ $0.10^{+0.04}_{-0.02}$ $33.2$ $5.13$
4 $3.7^{+1.5}_{-1.3}$ $5.6^{+1.8}_{-1.6}$ $2.41^{+0.02}_{-0.02}$ $0.10^{+0.04}_{-0.02}$ $33.3$ $4.88$
  • Table columns: (1) Co-addition method (SW or CW); (2) Dimensionless (in $\theta_{500}$ units) binning resolution; (3) Stacked catalog (2RXS, NVSS, or joint catalogs); (4) Effective VS model (planar or shell); (5) Effective number of free parameters; (6) and (7) Normalized, $n_{0}^{*}\equiv 10^{3}n_{0}$ X-ray and radio source densities in Eqs. (3.24) or (3.26); (8) Dimensionless shock radius; (9) Dimensionless signal width; (10) TS value; (11) TS-based significance. Numbers in bold are imposed constraints.
Appendix E Additional VS excess-source properties
Figures E.1E.6 demonstrate additional properties of the VS-excess sources, inferred in the same statistical method used in §5.
Figure E.1 shows that the black-body X-ray fit indicates similar, $L_{X}\simeq 10^{42\mbox{\scriptsize{--}}43}\mbox{ erg}\mbox{ s}^{-1}$ luminosities as those inferred from the power-law fit. Next, a comparison of normalized and non-normalized sizes is shown for X-ray (Fig. E.2) and radio (Fig. E.3) sources.
Figure E.1: Distribution of X-ray luminosity $\mathbb{P}=\log_{10}L_{X}$ according to power-law (left panel) and black-body (right) spectral fits. Notations are the same as in Fig. 11.
Figure E.2: Distributions of normalized (to $\theta_{500}$), dimensionless $\mathbb{P}=\tau_{R}$ (left panel) and non-normalized $\mathbb{P}=R$ (right) 2RXS source radii. Notations are the same as in Fig. 7.
Figure E.3: The same as Fig. E.2, but for the semi-major ($\mathbb{P}=a$, upper panels) and semi-minor ($\mathbb{P}=b$, lower panels) axes of radio sources. A variance $\sigma_{\rm smooth}^{2}(x)=(0.2x)^{2}$ is added to each scale $x$, for visibility.
Figure E.4 shows that extended radio sources may preferentially be oriented with their long axis perpendicular to the cluster radius, i.e. $|\varphi|\simeq\pi/2$, but this result is of low significance, based on only several low-declination (see Fig. 14) sources.
Figure E.4: Distribution of the angle $\mathbb{P}=|\varphi|$ between source elongation and cluster radius, as in Fig. 14, but for extended, $\theta_{a}>60{{}^{\prime\prime}}$ sources.
[…]
Figure E.5: Significance $S(\tau)$ profiles of SW-stacked NVSS (red up triangles), GMRT (orange diamonds), and VLSSr (blue down triangles) sources around the same clusters used for the NVSS sample. A low, $\Delta\tau=1/3$ resolution is used here, due to the smaller low-frequency catalogs.
Figure E.6: Same luminosity ${\mathbb{P}}=\log_{10}\nu L_{\nu}$ distribution as in Fig. 8, but for GMRT sources, around the same clusters as in the NVSS sample.
Appendix F No significant 2RXS-NVSS pair excess
In search of excess VS sources that radiate both in X-rays and in radio, we construct a catalog of 2RXS–NVSS pairs with an angular separation $<1^{\prime}$, chosen such that $<30\%$ of these pairs are likely to be coincidental projected overlaps. Stacking these pairs around our MCXC cluster sample does not produce a robust VS signal, as demonstrated in Fig. G.1. We conclude that the X-ray and radio excess VS signals are not dominated by the same individual objects.
Appendix G Excess sources are not identified AGN
We examine the possible association of excess VS sources with previously-identified AGN. For X-rays, 2RXS already provides associations with an AGN catalog [Veron_QSOs], whereas in radio, we cross-correlate NVSS with the same catalog using a proximity criterion of $<3^{\prime}$ angular separation. Almost all AGN associated with sources located in projection around an MCXC cluster are found to have redshifts inconsistent with the cluster. In addition, we find no significant difference in the results of stacking sources with vs. without an AGN association. Hence, we can rule out the possibility that a significant fraction of the sources comprising the excess VS signal is associated with these identified AGN.
[…]
Figure G.1: Excess significance $S(\tau)$ (panels a and b; see legends) and total number $N(\tau)$ (panel c; real pairs as disks with guidelines to the eye; expected field pairs as squares) profiles of 2RXS-NVSS pairs. Imposing the nominal flux cuts (solid green lines in panel c) leaves insufficient pairs. Relaxing $F_{\nu}^{*}$ (panel a and dashed blue in panel c) yields no signal. Relaxing $F_{X}^{*}$ (panel b and dot-dashed red in panel c) or both $F_{\nu}^{*}$ and $F_{X}^{*}$ (dotted black in panel c) leaves strong fluctuations, with a $\tau>2.5$ signal we cannot substantiate.
Appendix H VS-region sources
Tables LABEL:tab:VS_2RXS_list and LABEL:tab:VS_NVSS_list list respectively the 2RXS and NVSS sources found in the VS region, naming the cluster tentatively associated with each source. Note that statistically, only about $\sim 1/3$ of these sources are associated with the localized VS excess, the rest attributed to field (foreground or background) sources.
Table 4: 2RXS sources in the VS region
Source R.A. Dec Cluster Alt. name $z$
(1) (2) (3) (4) (5) (6)
2RXS J213734.6-720725 324.3927 -72.1233 MCXC J2133.4-7156 0.0559
2RXS J043118.7-620535 67.8281 -62.0940 MCXC J0431.4-6126 A3266 0.0589
2RXS J034914.6-544813 57.3109 -54.8043 MCXC J0352.3-5453 RBS 0485 0.0447
2RXS J034324.8-530041 55.8534 -53.0122 MCXC J0342.8-5338 A3158 0.0590
2RXS J034517.9-530544 56.3248 -53.0963 MCXC J0342.8-5338 A3158 0.0590
2RXS J033852.2-532427 54.7174 -53.4082 MCXC J0342.8-5338 A3158 0.0590
2RXS J034710.8-533449 56.7951 -53.5811 MCXC J0342.8-5338 A3158 0.0590
2RXS J034329.7-541259 55.8740 -54.2171 MCXC J0342.8-5338 A3158 0.0590
2RXS J224403.1-525611 341.0127 -52.9362 MCXC J2246.3-5243 A3911 0.0965
2RXS J032345.2-513226 50.9385 -51.5412 MCXC J0321.9-5119 A3120 0.0696
2RXS J032133.8-514003 50.3907 -51.6683 MCXC J0321.9-5119 A3120 0.0696
2RXS J024638.7-461700 41.6615 -46.2840 MCXC J0245.2-4627 A3047 0.0868
2RXS J034211.4-455724 55.5477 -45.9575 MCXC J0340.8-4542 RBS 0459 0.0698
2RXS J224909.0-453551 342.2874 -45.5975 MCXC J2250.8-4521 S1067 0.0511
2RXS J232040.0-424706 350.1665 -42.7850 MCXC J2319.1-4206 ACOS1111 0.0450
2RXS J225144.6-371317 342.9362 -37.2214 MCXC J2249.2-3727 S1065 0.0289
2RXS J225159.4-373447 342.9978 -37.5797 MCXC J2249.2-3727 S1065 0.0289
2RXS J053523.7-362736 83.8495 -36.4610 MCXC J0533.3-3619 S0535 0.0479
2RXS J235847.4-340950 359.6976 -34.1640 MCXC J2357.0-3445 A4059 0.0475
2RXS J000606.1-350150 1.5255 -35.0306 MCXC J0006.0-3443 A2721 0.1147
2RXS J002301.4-330721 5.7563 -33.1225 MCXC J0025.5-3302 S0041 0.0491
2RXS J052401.7-320833 81.0078 -32.1435 MCXC J0525.5-3135 A3341 0.0380
2RXS J040615.2-304956 61.5638 -30.8330 MCXC J0408.2-3053 A3223 0.0600
2RXS J001131.5-282046 2.8817 -28.3462 MCXC J0011.3-2851 A2734 0.0620
2RXS J001247.7-291301 3.1991 -29.2171 MCXC J0011.3-2851 A2734 0.0620
2RXS J005049.8-284533 12.7079 -28.7594 MCXC J0051.3-2831 A2829 0.1125
2RXS J024421.9-262500 41.0917 -26.4172 MCXC J0244.1-2611 0.1362
2RXS J001358.9-233820 3.4958 -23.6391 MCXC J0015.4-2350 A14 0.0645
2RXS J010013.3-215003 15.0559 -21.8344 MCXC J0102.7-2152 A0133 0.0569
2RXS J005002.6-212709 $1$21.4528 MCXC J0048.6-2114 A2824 0.0581
2RXS J215644.7-201430 329.1866 -20.2414 MCXC J2158.3-2006 A2401 0.0570
2RXS J045714.9-181757 74.3127 -18.3002 MCXC J0454.8-1806 CID 28 0.0335
2RXS J011459.1-153909 18.7467 -15.6530 MCXC J0116.1-1555 SCG 16 0.0448
2RXS J010832.7-155511 17.1369 -15.9202 MCXC J0108.9-1537 A0151S 0.0970
2RXS J010905.1-145729 17.2717 -14.9584 MCXC J0108.8-1524 A0151N 0.0533
2RXS J011029.5-151015 17.6233 -15.1712 MCXC J0108.8-1524 A0151N 0.0533
2RXS J005725.1-142056 14.3552 -14.3492 MCXC J0058.4-1425 A123 0.0957
2RXS J103914.9-090431 159.8131 -9.0757 MCXC J1039.7-0841 A1069 0.0650
2RXS J101848.3-062729 154.7021 -6.4585 MCXC J1020.4-0631 A0978 0.0540
2RXS J234648.0-014221 356.7006 -1.7059 MCXC J2347.4-0218 HCG 97 0.0223
2RXS J003606.6-021227 9.0281 -2.2078 MCXC J0034.6-0208 0.0812
2RXS J003545.5-020703 8.9403 -2.1178 MCXC J0034.2-0204 SH518 0.0822
2RXS J011731.8+002126 19.3833 0.3570 MCXC J0115.2+0019 A0168 0.0450
2RXS J011326.7+000035 18.3620 0.0095 MCXC J0115.2+0019 A0168 0.0450
2RXS J011703.9+000025 19.2671 0.0069 MCXC J0115.2+0019 A0168 0.0450
2RXS J213637.9+004203 324.1585 0.7016 MCXC J2137.1+0026 VMF98 202 0.0509
2RXS J013243.2+004109 23.1806 0.6856 MCXC J0131.7+0033 A208 0.0798
2RXS J013038.7+004006 22.6622 0.6683 MCXC J0131.7+0033 A208 0.0798
2RXS J010620.9+022017 16.5877 2.3380 MCXC J0108.1+0210 A0147 0.0436
2RXS J121929.8+034850 184.8748 3.8140 MCXC J1217.6+0339 ZW 1215.1+0400 0.0766
2RXS J122710.3+090646 186.7935 9.1131 MCXC J1227.4+0849 A1541 0.0896
2RXS J025806.4+093216 44.5277 9.5373 MCXC J0255.8+0918 IC 1867 0.0258
2RXS J110233.0+105722 165.6382 10.9561 MCXC J1100.8+1033 A1142 0.0353
2RXS J235542.9+114259 358.9297 11.7166 MCXC J2355.6+1120 A2675 0.0726
2RXS J224939.6+110016 342.4157 11.0048 MCXC J2250.0+1137 RX J2250.0+1137 0.0255
2RXS J114410.5+151430 176.0442 15.2417 MCXC J1145.3+1529 A1371 0.0670
2RXS J011459.8+152854 18.7499 15.4816 MCXC J0113.0+1531 A0160 0.0442
2RXS J145502.0+164416 223.7590 16.7386 MCXC J1454.4+1622 0.0456
2RXS J145502.0+164416 223.7590 16.7386 MCXC J1452.9+1641 A1983 0.0444
2RXS J145427.1+162242 223.6136 16.3790 MCXC J1452.9+1641 A1983 0.0444
2RXS J160624.0+181522 241.6005 18.2569 MCXC J1604.5+1743 A2151A 0.0370
2RXS J160726.2+175017 241.8596 17.8390 MCXC J1604.5+1743 A2151A 0.0370
2RXS J160624.0+181522 241.6005 18.2569 MCXC J1606.8+1746 A2151B 0.0321
2RXS J124142.2+181103 190.4265 18.1845 MCXC J1241.3+1834 A1589 0.0730
2RXS J232047.5+182931 350.1989 18.4924 MCXC J2318.5+1842 A2572 0.0403
2RXS J232047.5+182931 350.1989 18.4924 MCXC J2318.4+1843 A2572B 0.0389
2RXS J111124.7+212406 167.8532 21.4016 MCXC J1109.7+2145 A1177 0.0319
2RXS J112204.0+245206 170.5169 24.8684 MCXC J1122.3+2419 HCG 51 0.0258
2RXS J083659.6+243500 129.2488 24.5828 MCXC J0838.1+2506 CGCG120-014 0.0286
2RXS J142324.0+264159 215.8503 26.7003 MCXC J1423.1+2615 RX J1423.1+2616 0.0375
2RXS J134738.5+271451 206.9107 27.2479 MCXC J1348.8+2635 A1795 0.0622
2RXS J135006.7+271253 207.5284 27.2153 MCXC J1348.8+2635 A1795 0.0622
2RXS J134755.3+255544 206.9809 25.9292 MCXC J1348.8+2635 A1795 0.0622
2RXS J233319.1+271429 353.3308 27.2417 MCXC J2335.0+2722 A2622 0.0613
2RXS J152419.8+275742 231.0828 27.9625 MCXC J1522.4+2742 A2065 0.0723
2RXS J120914.3+280712 182.3100 28.1201 MCXC J1206.6+2811 NGC 4104 0.0283
2RXS J120745.5+273652 181.9399 27.6145 MCXC J1206.6+2811 NGC 4104 0.0283
2RXS J091750.4+341607 139.4603 34.2682 MCXC J0919.8+3345 A779 0.0230
2RXS J093429.7+333438 143.6241 33.5770 MCXC J0933.4+3403 UGC 05088 0.0269
2RXS J133712.0+344230 204.3001 34.7087 MCXC J1334.3+3441 NGC 5223 0.0240
2RXS J102045.2+381116 155.1885 38.1876 MCXC J1022.0+3830 RX J1022.1+3830 0.0491
2RXS J102128.8+380655 155.3702 38.1150 MCXC J1022.0+3830 RX J1022.1+3830 0.0491
2RXS J131156.8+385139 197.9870 38.8612 MCXC J1311.1+3913 A1691 0.0720
2RXS J120722.0+385213 181.8417 38.8705 MCXC J1205.2+3920 RX J1205.1+3920 0.0381
2RXS J171026.5+400332 257.6108 40.0598 MCXC J1711.0+3941 A2250 0.0647
2RXS J163031.2+410137 247.6304 41.0277 MCXC J1627.6+4055 A2197 0.0301
2RXS J162959.4+423008 247.4980 42.5030 MCXC J1627.3+4240 A2192 0.0317
2RXS J141348.9+440015 213.4539 44.0047 MCXC J1413.7+4339 A1885 0.0890
2RXS J141229.3+435534 213.1223 43.9267 MCXC J1413.7+4339 A1885 0.0890
2RXS J171124.1+434903 257.8508 43.8184 MCXC J1714.3+4341 NGC 6329 0.0276
2RXS J152115.3+485847 230.3139 48.9806 MCXC J1520.9+4840 A2064 0.1076
2RXS J113803.7+490500 174.5151 49.0835 MCXC J1134.8+4903 A1314 0.0341
2RXS J113408.7+483414 173.5360 48.5705 MCXC J1134.8+4903 A1314 0.0341
2RXS J161157.2+490138 242.9884 49.0280 MCXC J1614.1+4909 A2169 0.0579
2RXS J142404.6+493154 216.0190 49.5323 MCXC J1421.5+4933 ZW 1420.2+4952 0.0716
2RXS J142354.7+492828 215.9781 49.4749 MCXC J1421.5+4933 ZW 1420.2+4952 0.0716
2RXS J141945.7+491822 214.9405 49.3067 MCXC J1421.5+4933 ZW 1420.2+4952 0.0716
2RXS J102406.0+500302 156.0250 50.0504 MCXC J1022.5+5006 A980 0.1580
2RXS J100837.2+543735 152.1549 54.6262 MCXC J1010.2+5430 VMF98 84 0.0450
2RXS J105939.2+563211 164.9127 56.5363 MCXC J1058.4+5647 A1132 0.1369
2RXS J143027.2+565003 217.6129 56.8346 MCXC J1428.4+5652 A1925 0.1051
2RXS J162948.3+584616 247.4515 58.7720 MCXC J1629.7+5831 A2208 0.1329
2RXS J133520.4+584501 203.8343 58.7508 MCXC J1336.1+5912 A1767 0.0701
2RXS J170902.9+641730 257.2623 64.2925 MCXC J1712.7+6403 A2255 0.0809
2RXS J171640.0+641048 259.1670 64.1808 MCXC J1712.7+6403 A2255 0.0809
2RXS J171708.3+640146 259.2850 64.0304 MCXC J1712.7+6403 A2255 0.0809
2RXS J175341.7+654242 268.4244 65.7126 MCXC J1751.1+6531 NGC6505 0.0428
2RXS J174833.2+652351 267.1389 65.3985 MCXC J1751.1+6531 NGC6505 0.0428
2RXS J113146.0+660731 172.9404 66.1254 MCXC J1133.2+6622 A1302 0.1160
2RXS J174532.3+674811 266.3849 67.8039 MCXC J1742.7+6735 0.0420
2RXS J174603.5+672706 266.5151 67.4525 MCXC J1742.7+6735 0.0420
2RXS J175614.0+680708 269.0588 68.1198 MCXC J1755.7+6752 0.0833
2RXS J173039.0+681130 262.6628 68.1925 MCXC J1736.3+6803 ZW 1745.6+6703 0.0248
2RXS J174241.6+680013 265.6739 68.0045 MCXC J1736.3+6803 ZW 1745.6+6703 0.0248
2RXS J170234.2+681129 255.6427 68.1922 MCXC J1659.6+6826 0.0504
2RXS J170102.2+680513 255.2593 68.0878 MCXC J1659.6+6826 0.0504
2RXS J172207.7+694303 260.5323 69.7183 MCXC J1724.2+6956 0.0386
2RXS J172400.1+694026 261.0006 69.6749 MCXC J1724.2+6956 0.0386
2RXS J115509.4+734312 178.7870 73.7202 MCXC J1156.0+7325 A1412 0.0830
  • Table columns: (1) Source name; (2) Right ascension (deg); (3) Declination (deg); (4) Name of tentatively-associated cluster; (5) Alternative cluster name; (6) Cluster redshift. Note that each of the two sources, 2RXS J145502.0+164416 and 2RXS J232047.5+182931, appears twice, as it falls in the VS radii of two clusters. Such double counting affects field sources, too, and has a negligible effect on our results.
Table 5: NVSS sources in the VS region
Source R.A. Dec Cluster Alt. name $z$
(1) (2) (3) (4) (5) (6)
NVSS J000029-345224 0.1227 -34.8734 MCXC J2357.0-3445 A4059 0.0475
NVSS J011413-321800 18.5563 -32.3002 MCXC J0113.9-3145 S141A 0.0191
NVSS J011448-321951 18.7037 -32.3310 MCXC J0113.9-3145 S141A 0.0191
NVSS J011632-314001 19.1349 -31.6671 MCXC J0113.9-3145 S141A 0.0191
NVSS J004839-294718 12.1645 -29.7885 MCXC J0049.4-2931 S0084 0.1084
NVSS J005037-292128 12.6577 -29.3578 MCXC J0049.4-2931 S0084 0.1084
NVSS J000956-282923 2.4846 -28.4898 MCXC J0011.3-2851 A2734 0.0620
NVSS J001001-282940 2.5052 -28.4944 MCXC J0011.3-2851 A2734 0.0620
NVSS J054438-253222 86.1613 -25.5395 MCXC J0545.4-2556 A0548W 0.0424
NVSS J054916-260421 87.3176 -26.0727 MCXC J0548.6-2527 A0548E 0.0420
NVSS J055016-255502 87.5699 -25.9173 MCXC J0548.6-2527 A0548E 0.0420
NVSS J001403-233834 3.5152 -23.6430 MCXC J0015.4-2350 A14 0.0645
NVSS J234326-215422 355.8610 -21.9062 MCXC J2344.4-2153 A2655 0.1122
NVSS J220245-214333 330.6908 -21.7259 MCXC J2203.8-2130 0.0732
NVSS J055158-211948 87.9928 -21.3302 MCXC J0552.8-2103 A0550 0.0989
NVSS J082717-202622 126.8221 -20.4397 MCXC J0826.7-2007 S611 0.0876
NVSS J215707-202441 329.2812 -20.4116 MCXC J2158.3-2006 A2401 0.0570
NVSS J023743-193235 39.4324 -19.5432 MCXC J0236.6-1923 A0367 0.0907
NVSS J221145-122639 332.9387 -12.4442 MCXC J2210.3-1210 A2420 0.0846
NVSS J091541-114807 138.9248 -11.8022 MCXC J0918.1-1205 A0780 0.0539
NVSS J104108-114139 160.2867 -11.6944 MCXC J1041.5-1123 0.0839
NVSS J221519-090004 333.8319 -9.0011 MCXC J2216.2-0920 A2428 0.0825
NVSS J094446-082722 146.1937 -8.4562 MCXC J0945.4-0839 A0868 0.1535
NVSS J205703-075258 314.2643 -7.8830 MCXC J2058.2-0745 A2331 0.0793
NVSS J234515-035930 356.3143 -3.9918 MCXC J2344.2-0422 0.0786
NVSS J033834-022200 54.6449 -2.3668 MCXC J0340.6-0239 0.0352
NVSS J003438-014309 8.6595 -1.7192 MCXC J0034.2-0204 SH518 0.0822
NVSS J222347-020135 335.9484 -2.0265 MCXC J2223.8-0138 A2440 0.0906
NVSS J101444-000417 153.6845 -0.0715 MCXC J1013.7-0006 A0954 0.0927
NVSS J234057+000447 355.2402 0.0797 MCXC J2341.1+0018 0.1100
NVSS J011546+005204 18.9441 0.8679 MCXC J0115.2+0019 A0168 0.0450
NVSS J213638+004154 324.1607 0.6985 MCXC J2137.1+0026 VMF98 202 0.0509
NVSS J101135+005747 152.8969 0.9631 MCXC J1013.6-0054 A0957 0.0445
NVSS J091800+003450 139.5004 0.5808 MCXC J0920.0+0102 MKW 1S 0.0175
NVSS J092035+002330 140.1490 0.3918 MCXC J0920.0+0102 MKW 1S 0.0175
NVSS J092124+013834 140.3506 1.6428 MCXC J0920.0+0102 MKW 1S 0.0175
NVSS J023101+014958 37.7542 1.8330 MCXC J0231.9+0114 RCS145 0.0221
NVSS J102710+033927 156.7941 3.6577 MCXC J1028.3+0346 A1024 0.0730
NVSS J235254+060033 358.2277 6.0092 MCXC J2350.8+0609 A2665 0.0562
NVSS J003808+071223 9.5341 7.2066 MCXC J0040.0+0649 A76 0.0395
NVSS J082152+080430 125.4679 8.0751 MCXC J0821.0+0751 RX J0820.9+0751 0.1100
NVSS J122631+090037 186.6315 9.0105 MCXC J1227.4+0849 A1541 0.0896
NVSS J041528+103317 63.8691 10.5548 MCXC J0413.4+1028 A0478 0.0882
NVSS J114049+105754 175.2069 10.9651 MCXC J1141.2+1044 A1345 0.1090
NVSS J122901+120012 187.2546 12.0034 MCXC J1229.9+1147 A1552 0.0852
NVSS J122935+120548 187.3960 12.0968 MCXC J1229.9+1147 A1552 0.0852
NVSS J123056+115933 187.7336 11.9927 MCXC J1229.9+1147 A1552 0.0852
NVSS J025829+133328 44.6227 13.5579 MCXC J0257.8+1302 A0399 0.0722
NVSS J025831+133417 44.6329 13.5715 MCXC J0257.8+1302 A0399 0.0722
NVSS J025914+132735 44.8122 13.4598 MCXC J0257.8+1302 A0399 0.0722
NVSS J025634+133435 44.1457 13.5765 MCXC J0258.9+1334 A0401 0.0739
NVSS J232309+140444 350.7902 14.0789 MCXC J2324.3+1439 A2593 0.0428
NVSS J091204+161829 138.0167 16.3083 MCXC J0912.4+1556 A763 0.0851
NVSS J145602+162702 224.0084 16.4508 MCXC J1454.4+1622 0.0456
NVSS J145605+162652 224.0235 16.4480 MCXC J1454.4+1622 0.0456
NVSS J145608+162646 224.0339 16.4462 MCXC J1454.4+1622 0.0456
NVSS J123535+164732 188.8974 16.7924 MCXC J1236.4+1631 A1569 0.0780
NVSS J145420+162055 223.5846 16.3488 MCXC J1452.9+1641 A1983 0.0444
NVSS J145430+162250 223.6253 16.3806 MCXC J1452.9+1641 A1983 0.0444
NVSS J232609+171745 351.5393 17.2960 MCXC J2323.8+1648 A2589 0.0416
NVSS J160148+175401 240.4511 17.9004 MCXC J1604.5+1743 A2151A 0.0370
NVSS J160616+181459 241.5669 18.2498 MCXC J1604.5+1743 A2151A 0.0370
NVSS J160616+181459 241.5669 18.2498 MCXC J1606.8+1746 A2151B 0.0321
NVSS J075000+182311 117.5014 18.3866 MCXC J0748.1+1832 0.0400
NVSS J231822+191452 349.5954 19.2478 MCXC J2317.1+1841 A2572A 0.0422
NVSS J232046+182925 350.1948 18.4905 MCXC J2318.5+1842 A2572 0.0403
NVSS J232046+182925 350.1948 18.4905 MCXC J2318.4+1843 A2572B 0.0389
NVSS J003912+213405 9.8002 21.5681 MCXC J0039.6+2114 A0075 0.0619
NVSS J104702+221033 161.7617 22.1761 MCXC J1048.7+2214 A1100 0.0450
NVSS J115427+233804 178.6144 23.6346 MCXC J1155.3+2324 A1413 0.1427
NVSS J160212+241010 240.5525 24.1695 MCXC J1604.9+2355 AWM 4 0.0326
NVSS J115752+240748 179.4701 24.1302 MCXC J1156.9+2415 ZWCL4673 0.1392
NVSS J112145+245201 170.4414 24.8672 MCXC J1122.3+2419 HCG 51 0.0258
NVSS J112331+235047 170.8832 23.8465 MCXC J1122.3+2419 HCG 51 0.0258
NVSS J005122+242232 12.8430 24.3756 MCXC J0049.8+2426 A104 0.0815
NVSS J204130+244822 310.3752 24.8062 MCXC J2042.1+2426 0.1019
NVSS J204140+244924 310.4190 24.8234 MCXC J2042.1+2426 0.1019
NVSS J005655+263125 14.2296 26.5236 MCXC J0055.9+2622 A115 0.1971
NVSS J134612+261413 206.5509 26.2371 MCXC J1348.8+2635 A1795 0.0622
NVSS J134756+255812 206.9839 25.9702 MCXC J1348.8+2635 A1795 0.0622
NVSS J135011+271338 207.5472 27.2275 MCXC J1348.8+2635 A1795 0.0622
NVSS J172026+265456 260.1105 26.9157 MCXC J1720.1+2637 RX J1720.1+2638 0.1644
NVSS J005654+270909 14.2266 27.1527 MCXC J0058.9+2657 RX J0058.9+2657 0.0451
NVSS J111018+292016 167.5758 29.3379 MCXC J1110.7+2842 A1185 0.0314
NVSS J075248+294630 118.2031 29.7750 MCXC J0753.4+2921 A602 0.0621
NVSS J004204+294712 10.5179 29.7869 MCXC J0040.4+2933 A0077 0.0712
NVSS J154034+310239 235.1442 31.0442 MCXC J1539.8+3043 A2110 0.0980
NVSS J015715+315417 29.3137 31.9049 MCXC J0157.4+3213 A278 0.0894
NVSS J170011+323533 255.0491 32.5926 MCXC J1659.0+3229 0.0621
NVSS J170013+322549 255.0566 32.4305 MCXC J1659.0+3229 0.0621
NVSS J132218+325254 200.5753 32.8818 MCXC J1320.2+3308 NGC 5098 0.0362
NVSS J160047+325445 240.1991 32.9126 MCXC J1600.3+3313 A2145 0.0880
NVSS J101735+334031 154.3995 33.6754 MCXC J1016.3+3338 A961 0.1240
NVSS J174018+350549 265.0763 35.0969 MCXC J1740.5+3538 RX J1740.5+3539 0.0428
NVSS J174132+360724 265.3845 36.1235 MCXC J1740.5+3538 RX J1740.5+3539 0.0428
NVSS J105544+374601 163.9356 37.7671 MCXC J1057.7+3738 0.0353
NVSS J185527+374256 283.8652 37.7156 MCXC J1857.6+3800 0.0567
NVSS J015823+413739 29.5992 41.6276 MCXC J0157.1+4120 A276 0.0810
NVSS J084513+423918 131.3047 42.6553 MCXC J0844.9+4258 0.0541
NVSS J171123+434842 257.8475 43.8118 MCXC J1714.3+4341 NGC 6329 0.0276
NVSS J171202+431555 258.0089 43.2653 MCXC J1714.3+4341 NGC 6329 0.0276
NVSS J171205+431701 258.0234 43.2839 MCXC J1714.3+4341 NGC 6329 0.0276
NVSS J063932+474721 99.8868 47.7892 MCXC J0638.1+4747 ZWCL 0634.1+4750 0.1740
NVSS J102310+475146 155.7940 47.8628 MCXC J1025.0+4750 A1003 0.0520
NVSS J090650+500334 136.7102 50.0597 MCXC J0907.8+4936 VV 196 0.0352
NVSS J180947+492544 272.4500 49.4291 MCXC J1811.0+4954 ZWCL8338 0.0501
NVSS J184203+502404 280.5142 50.4012 MCXC J1843.6+5021 0.1158
NVSS J164934+535817 252.3950 53.9716 MCXC J1649.2+5325 ARP 330, SHK 016 0.0290
NVSS J160016+533945 240.0697 53.6625 MCXC J1601.3+5354 A2149 0.1068
NVSS J124538+551134 191.4106 55.1929 MCXC J1247.3+5500 A1616 0.0830
NVSS J134442+555313 206.1759 55.8870 MCXC J1343.7+5538 A1783 0.0766
NVSS J114558+552222 176.4947 55.3730 MCXC J1147.3+5544 A1377 0.0510
NVSS J142824+563611 217.1032 56.6032 MCXC J1428.4+5652 A1925 0.1051
NVSS J172338+562857 260.9112 56.4827 MCXC J1723.3+5658 NGC 6370 0.0272
NVSS J165438+583617 253.6584 58.6047 MCXC J1654.7+5854 A2239 0.0869
NVSS J175341+654238 268.4246 65.7108 MCXC J1751.1+6531 NGC6505 0.0428
NVSS J163540+655813 248.9171 65.9705 MCXC J1635.8+6612 A2218 0.1709
NVSS J113025+662605 172.6060 66.4349 MCXC J1133.2+6622 A1302 0.1160
NVSS J072754+665812 111.9751 66.9701 MCXC J0724.9+6658 A0578 0.0863
NVSS J062025+673640 95.1045 67.6111 MCXC J0618.6+6724 A0554 0.1127
NVSS J175049+680826 267.7083 68.1406 MCXC J1754.6+6803 ZW 1754.5+6807 0.0770
NVSS J173339+683557 263.4159 68.5994 MCXC J1736.3+6803 ZW 1745.6+6703 0.0248
NVSS J185609+680634 284.0397 68.1095 MCXC J1853.9+6822 A2312 0.0928
NVSS J181702+694125 274.2606 69.6904 MCXC J1814.2+6939 A2301 0.0863
NVSS J185101+732316 282.7568 73.3879 MCXC J1847.2+7320 A2310 0.1125
NVSS J115626+730650 179.1125 73.1141 MCXC J1156.0+7325 A1412 0.0830
NVSS J171206+774613 258.0265 77.7704 MCXC J1718.1+7801 A2271 0.0584
NVSS J165135+782853 252.8989 78.4815 MCXC J1703.8+7838 A2256 0.0581
  • Table columns: (1) Source name; (2) Right ascension (deg); (3) Declination (deg); (4) Name of tentatively-associated cluster; (5) Alternative cluster name; (6) Cluster redshift. Note that the source NVSS J160616+181459 appears twice, as it falls in the VS radius of two clusters.